### v2

# Phase Selection and Analysis for Multi-frequency Multi-user RIS Systems Employing Subsurfaces in Correlated Ricean and Rayleigh Environments

Abstract
Phase selection design for reconfigurable intelligent surfaces (RISs) is a significant research challenge, as a closed-form optimal solution for a multi-user (MU) system is believed to be intractable. While existing methods achieve strong near-optimal performance, they typically entail high computational complexity. In this work, we take a different approach and propose a practical method that achieves competitive performance while substantially reducing computational complexity. To do so, we consider a RIS divided into subsurfaces. Each subsurface is designed specifically for one user, who is served on their own frequency band. The other subsurfaces (those not designed for this user) provide additional uncontrolled scattering. We derive the exact closed-form expression for the mean signal-to-noise ratio (SNR) for the proposed subsurface design (SD) when all channels experience correlated Ricean fading. We simplify this to find the mean SNR for line-of-sight (LoS) channels and channels experiencing correlated Rayleigh fading. An iterative SD (ISD) process is proposed, where subsurfaces are designed sequentially, and the phases that are already set are used to enhance the design of the remaining subsurfaces. This is extended to a converged ISD (CISD), where the ISD process is repeated multiple times until the SNR increases by less than a specified tolerance. The ISD and CISD both provide a performance improvement over SD, which increases as the number of RIS elements increases. The SD is significantly simpler than the lowest complexity MU method we know of, and despite each user having less bandwidth, the SD outperforms the existing method in some key scenarios. The SD is more robust to strongly LoS channels and clustered users, as it does not rely on spatial multiplexing like other MU methods. Combined with the complexity reduction, this makes the SD an attractive phase selection method.
I Introduction
Reconfigurable intelligent surfaces (RISs) are the focus of considerable research interest for next generation mobile communications, due to their ability to manipulate the wireless channel. RIS can enhance several aspects of system performance, including signal-to-noise ratio (SNR), rank deficiency, and blockage avoidance [1]. A significant body of work on RIS now exists, including design, testbeds, and performance analysis [2]. Standards development is underway, with ETSI releasing its first report on RIS in 2023 [3].
Typical RIS implementations involve large numbers of elements in an array, where each element can be set to control the phase of the reflected signal. A key challenge is selecting the RIS element reflection coefficients to maximize the total SNR at the target receiver. Analytically optimal closed-form RIS designs for multi-user (MU) systems are not tractable, and many existing methods heavily rely on iterative optimization [4, 5, 6, 7, 8, 9, 10]. While these offer excellent performance, they involve significant computational overhead. This becomes particularly challenging when the RIS is passive and lacks onboard processing. Moreover, passive RIS cannot transmit pilots or process feedback, complicating channel estimation, particularly as the number of users increases. Therefore, reducing both the computational and channel estimation requirements is essential for the practical deployment of RIS-assisted MU systems.
I-A Review of Related Literature
The design of RIS phase selection methods has primarily focused on either iterative optimization methods, or the use of artificial intelligence (AI) and machine learning (ML) approaches, as analytically optimal closed-form RIS designs for MU systems are not tractable.
The foundational work in [4] presents a sum rate maximization method for MU RIS systems, where the RIS phase shift matrix and the base station (BS) precoding matrix are alternately optimized while considering non-uniform RIS reflection amplitudes. A similar problem was solved in [5] with imperfect channel state information (CSI).
In [6], the authors proposed a weighted downlink (DL)uplink (UL) sum rate optimization framework in which the DL and UL precoding matrices and the RIS phase shift matrix are alternately optimized. For a single-BS system, [7] proposes multiple joint optimization strategies over the RIS phases and BS beamforming vectors to maximize the SNR, and for a dual-BS system, the authors propose a joint optimization approach over the RIS phases and BS transmit powers that maximizes the mean signal-to-interference-plus-noise ratio (SINR). In [8], the transmit power is minimized by alternately optimizing the RIS phase shifts (in this case selected from a discrete set) and the BS beamforming vector under imperfect CSI.
The works in [9, 10, 11, 12] focus on lower complexity iterative methods. In [9], a method is proposed that selects the RIS phases to increase the sum rate by reducing the total mean-squared error (MSE) of the receiver rather, and in [10], the precoder, combiner and RIS phases are jointly optimized to minimize the MSE. The work in [11] proposes a low complexity approach to RIS channel estimation and the optimization of discrete RIS phases to maximize the sum rate, and in [12], a low complexity joint optimization method is detailed for a system involving two RISs, where the RIS phases and BS beamforming are jointly optimized to maximize sum rate.
The application of AI/ML to the RIS phase selection problem has allowed optimal phases to be found for more complex and larger systems. The work in [13] trains and tests a fully connected neural network to find the RIS phases that minimize the bit error rate (BER) and maximize the sum rate even with imperfect CSI. The authors of [14] use deep reinforcement learning to jointly optimize the RIS phases and power allocation coefficients to solve a multi-objective optimization problem, where power consumption is minimized and data rate is maximized. In [15], unsupervised learning is used to jointly optimize the BS beamforming vector and the RIS phases to maximize the SNR in a large-scale RIS system with thousands of elements.
All of the methods discussed above are designed to achieve optimal or near-optimal performance, and they deliver excellent results. However, even the low-complexity approaches within these [9, 10, 11, 12] rely on iterative procedures and non-trivial matrix operations and optimization steps. As a result, their computational burden can still be significant.
I-B Contributions
In this paper, we take a different direction to the works discussed in the previous section, and target useful performance gains with an extremely low computational complexity RIS phase selection method in a practical orthogonal frequency-division multiplexing (OFDM)-based system. While large-scale systems with multiple RISs are an important topic for research, the availability of exact analytical results for RIS systems remains limited. Thus, beginning with a simpler case allows general system insights to be gained through exact expressions and can provide a foundation for analyzing more complex systems in the future.
In [16], we proposed a low-complexity subsurface design (SD) where a RIS is divided into subsurfaces and one subsurface is designed to serve one user equipment (UE). A MU system employing OFDM is considered. OFDM systems are ubiquitous in 4G and 5G systems, and are expected to be so in 6G systems also [17]. Each user operates on a different frequency, which enables simple matched filtering (MF) to be used at the receiver and removes inter-user interference (IUI). The elements designed for a specific UE are selected using the optimal single-user (SU) phase selection method proposed in [18], and are optimal for that UE due to the lack of IUI. The elements designed for other UEs provide uncontrolled scattering for this UE. This may assist the UE, particularly when there is little natural uncontrolled scattering in the environment. As the elements are designed for a single UE, only the channels from that UE through one RIS subsurface need to be estimated. This leads to a reduction in the channel estimation requirement proportional to the number of UEs compared with designs where all UEs operate in the same frequency band. The use of MF ensures that the receiver processing is as simple as possible. The SD is not intended to compete with the most advanced iterative methods, but rather makes an important contribution to a critical gap in RIS research, namely simple low-complexity MU phase selection schemes that are viable for real-world implementation.
Motivated by the identified research gaps, we significantly extend the preliminary work in [16] by broadening the analysis conducted to a much more general scenario, and proposing multiple extensions to the SD that offer further performance gains. The main contributions of this work are as follows:
  • In our preliminary work in [16], we only considered the performance of the SD for a line-of-sight (LoS) RIS-BS channel and correlated Rayleigh UE-BS and UE-RIS channels. In reality, it is likely that all channels will have both LoS and non-line-of-sight (NLoS) components. Therefore, we now significantly extend our work by investigating the performance of this simple and practical method for more general and realistic correlated Ricean fading channels. This versatile analytical tool can be simplified to more specific combinations of LoS and NLoS channels as desired.
  • We also extend the subsurface design to a low-complexity iterative subsurface design (ISD). The SD assumes that each subsurface is designed independently with no knowledge of phases set for other users. The ISD sets the subsurfaces sequentially, utilizing the knowledge of all phases set for other users to improve the design. The ISD performs this process once, but it can be performed repeatedly until convergence and the total SNR stops increasing. This is referred to as the converged ISD (CISD). The ISD and CISD offer performance improvements while still benefiting from low computational complexity and a simple MF receiver.
  • Using analysis and simulations, we investigate the performance of the SD, ISD and CISD. The mean SNR of each method is compared for a range of RIS dimensions. The mean sum rate supported by the SD is compared to a lower-bound benchmark equivalent to no channel knowledge (e.g. randomly selected RIS phases) and the MU method in [9]. This was selected for comparison as it is the lowest complexity existing full MU design that we are aware of, and while other methods may offer better performance, it is highly probable that they will require considerably higher system complexity. We thoroughly investigate the difference between the SD and MU method for a wide range of system parameters, including RIS size, correlation between BS and RIS elements, Ricean K-factor, UE location and UE clustering, and highlight important scenarios where the SD outperforms the MU method.
Notation: Upper and lower boldface letters represent matrices and vectors, respectively. We define vector $\mathbf{v}=[\mathbf{v}_{1}^{T},\dots,\mathbf{v}_{K}^{T}]^{T}$, where $\mathbf{v}_{k}$ is the $k$-th block of vector $\mathbf{v}$ and $\mathbf{v}_{k,i}$ is the $i$-th element of the $k$-th block. $\mathbf{M}$ is a matrix comprising of submatrices, or blocks. $\mathbf{M}_{r,s,i,j}$ is the $(i,j)$-th element of the $r$-th block across and $s$-th block down. Matrices are indexed this way except where explicitly defined otherwise. $\mathbf{M}^{(k)}$ denotes a matrix relating to user $k$. $\mathbb{C}$ is the set of complex numbers. $\Re[\cdot]$ is the real operator. $\mathcal{CN}(\boldsymbol{\mu},\mathbf{Q})$ represents a complex Gaussian distribution with mean $\boldsymbol{\mu}$ and covariance matrix $\mathbf{Q}$. $\mathbb{E}[\cdot]$ represents statistical expectation. $(\cdot)^{T}$, $(\cdot)^{*}$ and $(\cdot)^{\dagger}$ are the transpose, conjugate and Hermitian transpose operators, respectively. The angle of a vector, $\mathbf{x}$, of length $N$ is denoted as $\angle\mathbf{x}=[\angle{x}_{1},\dots,\angle{x}_{N}]^{T}$ and the exponent as $e^{\mathbf{x}}=[e^{{x}_{1}},\dots,e^{{x}_{N}}]^{T}$. $||\cdot||$ is the Euclidean norm. $_{1}F_{1}(a,b;z)$ is the confluent hypergeometric function and $_{2}F_{1}(a,b,c;z)$ is the Gaussian hypergeometric function. $L_{\nu}(\cdot)$ denotes a Laguerre function of non-integer degree $\nu$. $\Gamma(\cdot)$ is the complete gamma function. $f^{*}(r,\theta)$ is the probability density function (PDF) of a complex random variable $X=r\,\mathrm{e}^{j\theta}$, and $f^{*}(r_{1},r_{2},\theta_{1},\theta_{2})$ is the joint PDF of complex random variables $X_{1}=r_{1}\mathrm{e}^{j\theta_{1}}$ and $X_{2}=r_{2}\mathrm{e}^{j\theta_{2}}$. $J_{n}(\cdot)$ is the $n$-th order Bessel function of the first kind and $I_{n}(\cdot)$ is the $n$-th order modified Bessel function of the first kind.
II System Model
We consider the UL RIS-aided system shown in Fig. 1, consisting of a RIS panel with $N$ elements, a BS with $M$ antennas, and $K$ single-antenna users located in the vicinity of the BS and RIS. A RIS link (UE-RIS-BS) and a direct link (UE-BS) connect the UE to the BS. The system bandwidth, $B$, is split into $K$ bands of $\frac{B}{K}$ Hz, with one user per band. The phase response of the RIS, and therefore the phase shift in the reflected signal, is fixed across the frequency bands. The RIS elements are grouped into “subsurfaces” and each subsurface is designed for a different user.
Figure 1: System model showing UL channels for UE 1 in band 1.
For UE $k$ in band $k$, let $\mathbf{h}^{(k)}_{\mathrm{d}}\!\in\mathbb{C}^{M\times 1}$, $\mathbf{h}^{(k)}_{\mathrm{ur}}\in\mathbb{C}^{N\times 1}$ and $\mathbf{H}^{(k)}_{\mathrm{rb}}\in\mathbb{C}^{M\times N}$ be the direct, UE-RIS and RIS-BS channels, respectively. $\mathbf{\Phi}\in\mathbb{C}^{N\times N}$ is a diagonal matrix of reflection coefficients for the RIS which can be given in block diagonal form, where the $k$-th block is designed to enhance the channel for user $k$. $\mathbf{\Phi}=\mathrm{diag}\{\mathbf{\Phi}_{1},\dots,\mathbf{\Phi}_{K}\}$, where $\mathbf{\Phi}_{k}=\mathrm{diag}(e^{j\phi_{k,1}},\dots,e^{j\phi_{k,N_{k}}})$. $N_{k}=N/K$ is the number of RIS elements chosen to support UE $k$ and $N=\sum^{K}_{k=1}N_{k}$. We consider an equal number of elements per UE, and non-equally divided subsurfaces are a potential topic for future work. Elements serving one UE are co-located, as this was demonstrated to be the best configuration in [16]. The received signal in band $k$ at the BS is
$\mathbf{r}_{k}=(\mathbf{h}^{(k)}_{\mathrm{d}}+\mathbf{H}^{(k)}_{\mathrm{rb}}\mathbf{\Phi h}_{\mathrm{ur}}^{(k)})s_{k}+\mathbf{n}_{k},$ (1)
where $s_{k}$ is the signal being sent from UE $k$, $\mathbb{E}[|s_{k}|^{2}]=E_{s}$, and $\mathbf{n}_{k}\sim\mathbb{C}\mathcal{N}(0,\sigma^{2}\mathbf{I})$ is additive white Gaussian noise (AWGN). Using the block diagonal form of $\mathbf{\Phi}$, (1) becomes
$\mathbf{r}_{k}=(\mathbf{h}^{(k)}_{\mathrm{d}}+\mathbf{H}^{(k)}_{\mathrm{rb},k}\mathbf{\Phi}_{k}\mathbf{h}^{(k)}_{\mathrm{ur},k})s_{k}\\ +\left(\sum_{s\neq k}^{K}\mathbf{H}^{(k)}_{\mathrm{rb},s}\mathbf{\Phi}_{s}\mathbf{h}^{(k)}_{\mathrm{ur},s}\right)s_{k}+\mathbf{n}_{k},$ (2)
where $\mathbf{H}^{(k)}_{\mathrm{rb}}=\left[\mathbf{H}^{(k)}_{\mathrm{rb},1},\dots,\mathbf{H}^{(k)}_{\mathrm{rb},K}\right$], $\mathbf{H}^{(k)}_{\mathrm{rb},i}\in\mathbb{C}^{M\times N_{i}}$, $\mathbf{h}^{(k)}_{\mathrm{ur}}=\left[\mathbf{h}^{(k)T}_{\mathrm{ur},1},\dots,\mathbf{h}^{(k)T}_{\mathrm{ur},K}\right]^{T}$ and $\mathbf{h}^{(k)}_{\mathrm{ur},i}\in\mathbb{C}^{N_{i}\times 1}$.
II-A Channel Model
In this work, we consider correlated Ricean channels consisting of a rank-1 LoS and correlated Rayleigh component for all links. The LoS components are comprised of steering vectors and the scattered components adopt the Kronecker correlation model. Therefore, the channels for UE $k$ are
$\mathbf{h}^{(k)}_{\mathrm{d}}$ $=\sqrt{\beta^{(k)}_{\mathrm{d}}}\left(\eta^{(k)}_{\mathrm{d}}\tilde{\mathbf{h}}^{(k,\mathrm{LoS})}_{\mathrm{d}}+\zeta^{(k)}_{\mathrm{d}}\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{SC})}\right),$ (3)
$\mathbf{H}^{(k)}_{\mathrm{rb}}$ $=\sqrt{\beta^{(k)}_{\mathrm{rb}}}\left(\eta^{(k)}_{\mathrm{rb}}\tilde{\mathbf{H}}_{\mathrm{rb}}^{(k,\mathrm{LoS})}+\zeta^{(k)}_{\mathrm{rb}}\tilde{\mathbf{H}}^{(k,\mathrm{SC})}_{\mathrm{rb}}\right)$, (4)
$\mathbf{h}^{(k)}_{\mathrm{ur}}$ $=\sqrt{\beta^{(k)}_{\mathrm{ur}}}\left(\eta^{(k)}_{\mathrm{ur}}\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{LoS})}+\zeta^{(k)}_{\mathrm{ur}}\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{SC})}\right),$ (5)
with
[…]
where $\beta^{(k)}_{\mathrm{d}}$, $\beta^{(k)}_{\mathrm{rb}}$ and $\beta^{(k)}_{\mathrm{ur}}$ are the channel gains, $\kappa^{(k)}_{\mathrm{d}}$, $\kappa^{(k)}_{\mathrm{rb}}$ and $\kappa^{(k)}_{\mathrm{ur}}$ are the Ricean K-factors, $\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{LoS})}$, $\tilde{\mathbf{H}}_{\mathrm{rb}}^{(k,\mathrm{LoS})}$ and $\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{LoS})}$ are the LoS channel components for the $k$-th UE and $\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{SC})}$, $\tilde{\mathbf{H}}_{\mathrm{rb}}^{(k,\mathrm{SC})}$ and $\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{SC})}$ are the scattered channel (SC) components for the $k$-th UE. Here,
$\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{LoS})}$ $=\mathbf{a}^{(k)}_{\mathrm{d}}$, $\qquad\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{SC})}$ $=\mathbf{R}^{(k)1/2}_{\mathrm{d}}\mathbf{u}^{(k)}_{\mathrm{d}}$,
$\tilde{\mathbf{H}}_{\mathrm{rb}}^{(k,\mathrm{LoS})}$ $=\mathbf{a}^{(k)}_{\mathrm{b}}\mathbf{a}^{(k)\dagger}_{\mathrm{r}}$, $\qquad\tilde{\mathbf{H}}_{\mathrm{rb}}^{(k,\mathrm{SC})}$ $=\mathbf{R}_{\mathrm{b}}^{(k)1/2}\mathbf{U}_{\mathrm{rb}}^{(k)}\mathbf{R}_{\mathrm{r}}^{(k)1/2},$
$\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{LoS})}$ $=\mathbf{a}_{\mathrm{ur}}^{(k)},$ $\qquad\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{SC})}$ $=\mathbf{R}_{\mathrm{ur}}^{(k){1/2}}\mathbf{u}^{(k)}_{\mathrm{ur}}$,
where $\mathbf{a}^{(k)}_{\mathrm{d}}$ is the steering vector of the LoS ray for the direct link at the BS, $\mathbf{a}^{(k)}_{\mathrm{b}}$ and $\mathbf{a}^{(k)}_{\mathrm{r}}$ are steering vectors for the LoS ray for the RIS-BS link at the BS and RIS, respectively, and $\mathbf{a}^{(k)}_{\mathrm{ur}}$ is the steering vector of the LoS ray for the UE-RIS link at the RIS. $\mathbf{R}^{(k)}_{\mathrm{b}}$, $\mathbf{R}^{(k)}_{\mathrm{r}}$, $\mathbf{R}^{(k)}_{\mathrm{d}}$ and $\mathbf{R}^{(k)}_{\mathrm{ur}}$ are the correlation matrices for the RIS-BS link at the BS and RIS ends, the UE-BS link and UE-RIS link, respectively, and $\mathbf{u}^{(k)}_{\mathrm{d}},\mathbf{U}^{(k)}_{\mathrm{rb}}$ and $\mathbf{u}^{(k)}_{\mathrm{ur}}$ are all matrices and vectors containing independent and identically distributed $\mathbb{C}\mathcal{N}(0,1)$ entries.
This work focuses on strongly LoS $\mathbf{H}_{\mathrm{rb}}$ channels, where one ray is dominant but a scattered component is also present. This is a common and reasonable assumption for RIS systems, as the RIS can be located to exploit the LoS channel as seen in [19, 20, 21]. We also assume that reliable CSI is available for each individual link. Prior work has demonstrated that the UE-RIS and RIS-BS channels can be estimated separately, as in [22, 23]. Hence, we anticipate that analogous methods could be applied here.
III Phase Selection Methods
Partitioning the $N$ RIS elements into $K$ subsurfaces each designed independently for a specific UE enables low-complexity and efficient RIS phase selection. This section outlines three methods that utilise this concept: a simple approach that uses the SU optimal phase selection method for when $\mathbf{H}_{\mathrm{rb}}$ is rank-1, a sub-optimal extension of this for when $\mathbf{H}_{\mathrm{rb}}$ has a scattered component, and a higher-performing but more complex iterative version.
III-A LoS Phase Selection Method
As we first proposed in [16], the phases of a subsurface designed for the $k$-th UE can be set according to the SU method set out in [18], so that
$\mathbf{\Phi}_{k}=\nu_{k}\,\mathrm{diag}\left(e^{j\left(\angle\mathbf{a}^{(k)}_{\mathrm{r},k}-\angle\mathbf{h}^{(k)}_{\mathrm{ur},k}\right)}\right)$, (12)
where
$\nu_{k}=\frac{\mathbf{a}^{(k)\dagger}_{\mathrm{b}}\mathbf{h}^{(k)}_{\mathrm{d}}}{|\mathbf{a}^{(k)\dagger}_{\mathrm{b}}\mathbf{h}^{(k)}_{\mathrm{d}}|}$. (13)
The phases set for each UE are optimal for that UE when $\mathbf{H}_{\mathrm{rb}}^{(k)}$ is purely LoS, and close to optimal when $\mathbf{H}_{\mathrm{rb}}^{(k)}$ has a strong LoS component. This method leads to a $K$-fold reduction in the required CSI compared to typical RIS designs, due to only requiring the estimation of $N_{k}$ rather than $N$ channel elements per user. There are no complex matrix operations required and a simple MF receiver can be used. These advantages make this approach a very attractive option for implementation.
III-B Extended Subsurface Design for NLoS Systems
The design in (12) and (13) works well for strongly LoS channels. However, if there is a strong scattered component, a different approach provides better results. As there is no optimal closed form SU phase selection method when $\mathbf{H}^{(k)}_{\mathrm{rb}}$ is not rank-1, we propose the following adaptation to the method outlined in Sec. III-A, referred to as the extended subsurface design (ESD). As we proposed in [16], we take the singular value decomposition (SVD) of $\mathbf{H}^{(k)}_{\mathrm{rb},k}$, so that $\mathbf{H}^{(k)}_{\mathrm{rb},k}=\mathbf{UDV^{\dagger}}$, where $\mathbf{D}=\mathrm{diag}(D_{11},D_{22},\dots,D_{mm})$, $m=\min(M,N_{k})$ and the singular values are $D_{11}\geq D_{22}\geq...\geq D_{mm}$. The strongest rank-1 component of $\mathbf{H}^{(k)}_{\mathrm{rb},k}$ is $D_{11}\mathbf{u}_{k}\mathbf{v}_{k}^{\dagger}$, where $\mathbf{u}_{k}$ and $\mathbf{v}_{k}$ are the leading left and right singular vectors, respectively. This has the same structure as a pure LoS channel, so the same design can be used by considering just the strongest component, giving,
$\mathbf{\Phi}_{\mathrm{k}}=\omega_{k}\,\mathrm{diag}\left(e^{j\left(\mathbf{\angle v}_{k}-\mathbf{\angle h}^{(k)}_{\mathrm{ur},k}\right)}\right)$, (14)
where
$\omega_{k}=\frac{|\mathbf{h}_{\mathrm{ru},k}^{(k)}|^{T}\mathrm{diag}(e^{-j\angle\mathbf{v}_{k}})\mathbf{H}_{\mathrm{br},k}^{(k)\dagger}\mathbf{h}_{\mathrm{d}}^{(k)}}{\left||\mathbf{h}_{\mathrm{ru},k}^{(k)}|^{T}\mathrm{diag}(e^{-j\angle\mathbf{v}_{k}})\mathbf{H}_{\mathrm{br},k}^{(k)\dagger}\mathbf{h}_{\mathrm{d}}^{(k)}\right|}.$ (15)
III-C Iterative Phase Alignment
The subsurface methods outlined in Secs. III-A and III-B assume that each subsurface is designed independently, with no knowledge of phases that have been set for other users. Since the reflections produced by the other subsurfaces manifest as random scattering for a given user, setting the subsurfaces sequentially allows this randomness to be incorporated into the design and used to inform the configuration of the remaining subsurfaces.
The methods in the previous sections select phases to maximize the norm of the subsurface’s RIS link, $\mathbf{H}^{(k)}_{\mathrm{rb},k}\mathbf{\Phi}_{k}\mathbf{h}^{(k)}_{\mathrm{ur},k}$, and then rotate these phases to align with the non-RIS-controlled direct link, $\mathbf{h}^{(k)}_{\mathrm{d}}$. The ISD instead involves selecting phases to maximize the RIS link, and then aligning these with a combination of the direct path and the RIS paths that have already been set for other UEs. This is done by including this information into $\nu_{k}$ (LoS case) or $\omega_{k}$ (NLoS case). The subsurfaces are set from the weakest UE-RIS channel to the strongest, so that the most impactful channels can benefit from the most information. This process is detailed in Algorithm 1, where index()($\cdot$) refers to a function that returns the index of elements in a list and sort()($\cdot$) refers to a function that reorders elements of a list from smallest to largest.
This iterative method improves the mean SNR as shown in Sec. V-A at the cost of a significant increase in CSI requirements at the BS. $\mathbf{h}_{\mathrm{ur}}^{(k)}$ contains $N$ channel elements, with $\frac{N}{K}$ corresponding to those impacted by a specific subsurface. The iterative method requires the computation of the magnitude of a user’s UE-RIS link to rank channels based on their impact, so knowledge of the amplitude of all $N$ channel elements is required. In contrast, the methods in Secs. III-A and III-B only consider the impact on a UE from each subsurface, so only the subsurface specific $\frac{N}{K}$ elements are required. The benefits of the iterative method vary with channel conditions such as correlation and link power, as will be shown in Sec. V-A.
Algorithm 1 Iterative Phase Alignment Method
[…]
for $k=1$ to $K$ do
  $\mathit{index\_k}=\mathit{user\_order}(k)$
  $\mathit{fixed}=0$
  for $l=1$ to k1k-1 do
   $\mathit{index\_l}=\mathit{user\_order}(l)$
   $\mathit{fixed}=\mathit{fixed}+\mathbf{H}^{(\mathit{index\_k})}_{\mathrm{rb},\mathit{index\_l}}\mathbf{\Phi}_{\mathit{index\_l}}\mathbf{h}^{(\mathit{index\_k})}_{\mathrm{ur},\mathit{index\_l}}$
  end for
  $\mathit{direct}=\mathbf{h}^{(\mathit{index\_k})}_{\mathrm{d}}+\mathit{fixed}$
  $\nu_{\mathit{index\_k}}=\dfrac{\mathbf{a}_{\mathrm{b}}^{(\mathit{index\_k})\dagger}\times\mathit{direct}}{|\mathbf{a}_{\mathrm{b}}^{(\mathit{index\_k})\dagger}\times\mathit{direct}|}$
  $\mathbf{\Phi}_{\mathrm{\mathit{index\_k}}}=\nu_{\mathit{index\_k}}\,\mathrm{diag}\left(\mathrm{e}^{j\left(\angle\mathbf{a}^{(\mathit{index\_k})}_{\mathrm{r}}-\angle\mathbf{h}^{(\mathit{index\_k})}_{\mathrm{ur},\mathit{index\_l}}\right)}\right)$
end for
IV Analysis
In this section, we focus on an analysis of the mean SNR as prior work [18, 16, 24] has shown it to be tractable and to deliver system insights. The accuracy of this rate bound was investigated for the SD in our work in [16], and it was found to be tight for all values. In particular, we derive the mean SNR for a correlated Ricean environment, where $\mathbf{h}_{\mathrm{d}}$, $\mathbf{H}_{\mathrm{rb}}$ and $\mathbf{h}_{\mathrm{ur}}$ all consist of a rank-1 LoS and a scattered component. We consider the important case where $\mathbf{H}_{\mathrm{rb}}$ has a strong, but not pure, LoS component. This leads to a stable, slow-moving RIS-BS channel, allowing the SD to be used with close to optimal performance for a given subsurface [25]. Unfortunately, as far as we know, a closed-form result for the mean SNR is not available using the ESD. However, the SD remains highly accurate in the considered scenario. This enables the derivation of closed-form results, which are particularly valuable for understanding how system behavior arises from underlying mathematical structure. Therefore, $\mathbf{\Phi}$ in (12) is used in this section. This general case is then simplified to two special cases where $\mathbf{h}_{\mathrm{d}}$ and $\mathbf{h}_{\mathrm{ur}}$ are correlated Rayleigh channels; in the first $\mathbf{H}_{\mathrm{rb}}$ is correlated Ricean, and in the second it is a rank-1 LoS channel. The received signal $\mathbf{r}_{k}$ in (2) can be rewritten as
$\mathbf{r}_{k}=\left(\mathbf{h}^{(k)}_{\mathrm{d}}+\mathbf{f}_{k}+\mathbf{g}_{k}\right)s_{k}+\mathbf{n}_{k},$ (16)
where $\mathbf{f}_{k}=\mathbf{H}^{(k)}_{\mathrm{rb},k}\mathbf{\Phi}_{k}\mathbf{h}^{(k)}_{\mathrm{ur},k}$ and $\mathbf{g}_{k}=\sum_{s\neq k}^{K}\mathbf{H}^{(k)}_{\mathrm{rb},s}\mathbf{\Phi}_{s}\mathbf{h}^{(k)}_{\mathrm{ur},s}$. Thus, the SNR is given by
$\mathrm{SNR}_{k}=\frac{E_{s}}{\sigma^{2}}\Big[\mathbf{h}^{(k)\dagger}_{\mathrm{d}}\mathbf{h}^{(k)}_{\mathrm{d}}+2\Re(\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{f}_{k})+2\Re(\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{g}_{k})\\ +2\Re(\mathbf{f}_{k}^{\dagger}\mathbf{g}_{k})+\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}+\mathbf{g}_{k}^{\dagger}\mathbf{g}_{k}\Big].$ (17)
IV-A Mean SNR for Correlated Ricean Systems
We now consider the mean value of each term in (17) to calculate $\mathbb{E}[\mathrm{SNR}_{k}]$ for correlated Ricean systems. From [25],
$\mathbb{E}[\mathbf{h}^{(k)\dagger}_{\mathrm{d}}\mathbf{h}^{(k)}_{\mathrm{d}}]=M\beta_{\mathrm{d}}^{(k)},$ (18)
and
[…]
Remark 1
$\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{h}^{(k)}_{\mathrm{d}}$] represents the gain of the direct channel, and $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{f}_{\mathrm{k}}]$ is the cross product of the direct channel and channel through the subsurface designed for UE $k$. Higher element spacing at the BS increases the SNR, due to the $\mathbf{R}_{\mathrm{d}}\mathbf{a}_{\mathrm{b}}||$ component of (19). This term reduces if BS elements are placed closely together, as in [18].
All remaining terms are derived in Appendix A and their results are listed below.
$\mathbb{E}[\mathbf{h}^{(k)\dagger}_{\mathrm{d}}\mathbf{g}_{\mathrm{k}}]=\sqrt{\beta^{(k)}_{\mathrm{d}}\beta^{(k)}_{\mathrm{rb}}\beta^{(k)}_{\mathrm{ur}}}\eta^{(k)}_{\mathrm{d}}\eta^{(k)}_{\mathrm{rb}}\eta^{(k)}_{\mathrm{ur}}\mathbf{a}^{(k)\dagger}_{\mathrm{d}}\mathbf{a}_{\mathrm{b}}^{(k)}\\ \times\sum_{s\neq k}\mathbf{a}_{\mathrm{r},s}^{(k)\dagger}\mathbf{C}(s)\mathbf{a}_{\mathrm{ur},s}^{(k)},$ (20)
where
$\mathbf{C}(x)\!=\!\frac{\pi\sqrt{\!\kappa_{\mathrm{d}}^{(x)}\!\kappa_{\mathrm{ur}}^{(x)}}\mathbf{a}_{\mathrm{b}}^{(x)\dagger}\mathbf{a}_{\mathrm{d}}^{(x)}\!}{4\sqrt{\mathbf{a}_{\mathrm{b}}^{(x)\dagger}\mathbf{R}_{\mathrm{d}}^{(x)}\mathbf{a}_{\mathrm{b}}^{(x)}}}\mathrm{diag}\!\left(\!\mathrm{e}^{j\left(\angle\mathbf{a}_{\mathrm{r},x}^{(x)}-\angle\mathbf{a}_{\mathrm{ur},x}^{(x)}\right)}\!\right)\\ \times{}_{1}F_{1}\!\!\left(\!\frac{1}{2},2,-\frac{|\mathbf{a}^{(x)\dagger}_{\mathrm{b}}\mathbf{a}^{(x)}_{\mathrm{d}}|^{2}\kappa^{(x)}_{\mathrm{d}}}{\mathbf{a}_{\mathrm{b}}^{(x)\dagger}\mathbf{R}_{\mathrm{d}}^{(x)}\mathbf{a}_{\mathrm{b}}^{(x)}}\!\right)\!{}_{1}F_{1}\!\!\left(\!\frac{1}{2},2,-\kappa_{\mathrm{ur}}^{(x)}\!\right).$ (21)
Remark 2
$\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{g}_{\mathrm{k}}]$ represents the cross products of the direct channel and the channels through subsurfaces designed for other UEs. The complexity of the summation term in (20) makes it difficult to interpret in detail, but the size of the term is increased, as expected, by alignment of the direct and RIS paths. Note that the summation of a large number of phases in $\mathbf{a}^{(k)}_{\mathrm{r},s}$ and $\mathbf{a}^{(k)}_{\mathrm{ur},s}$ is likely to cause significant cancellation, so this is not a dominant term.
$\mathbb{E}[\mathbf{f}_{\mathrm{k}}^{\dagger}\mathbf{g}_{\mathrm{k}}]=\frac{M\pi\beta_{\mathrm{rb}}^{(k)}\beta_{\mathrm{ur}}^{(k)}\sqrt{\kappa_{\mathrm{d}}^{(k)}\kappa_{\mathrm{ur}}^{(k)}}\mathbf{a}_{\mathrm{d}}^{(k)\dagger}\mathbf{a}_{\mathrm{b}}^{(k)}\mathrm{e}^{-\kappa_{\mathrm{ur}}^{(k)}}}{4\!\left(1\!+\!\kappa_{\mathrm{rb}}^{(k)}\right)\!\!\left(1\!+\!\kappa_{\mathrm{ur}}^{(k)}\right)\!\sqrt{\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{R}_{\mathrm{d}}^{(k)}\mathbf{a}_{\mathrm{b}}^{(k)}}}$
$\times\!{}_{1}F_{1}\!\left(\!\frac{1}{2},2,-\frac{|\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{a}_{\mathrm{d}}^{(k)}|^{2}\kappa_{\mathrm{d}}^{(k)}}{\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{R}_{\mathrm{d}}^{(k)}\mathbf{a}_{\mathrm{b}}^{(k)}}\!\right)\!\sum_{s\neq k}\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}\bigg[\!\Big(\mathbf{a}^{(k)}_{\mathrm{ur},s,j}\!-\!\rho^{(k)}_{ks}$
$\times\mathbf{a}_{\mathrm{ur},k,i}^{(k)}\Big){}_{1}F_{1}\!\!\left(\!\frac{3}{2},1,\kappa_{\mathrm{ur}}^{(k)}\!\right)\!+\!\frac{3}{2}\rho^{(k)}_{ks}{}_{1}F_{1}\!\!\left(\!\frac{5}{2},2,\kappa_{\mathrm{ur}}^{(k)}\!\right)\!\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}}\!\bigg]$
$\times\!\left(\!\kappa_{\mathrm{rb}}^{(k)}\mathbf{a}^{(k)}_{\mathrm{r},k,i}\mathbf{a}^{(k)*}_{\mathrm{r},s,j}\!+\!r_{ks}^{(k)}\right)\!(\mathbf{C}(s))_{j,j}\,\mathrm{e}^{-j\angle\mathbf{a}_{\mathrm{r},k,i}^{(k)}},$ (22)
where $\mathbf{C}(x)$ is given in (21), and, dropping subscripts for readability, $\rho^{(k)}_{ks}=\mathbf{R}^{(k)}_{\mathrm{ur},k,s,i,j}$ and $r^{(k)}_{ks}=\mathbf{R}^{(k)}_{\mathrm{r},k,s,i,j}$.
Remark 3
$\mathbb{E}[\mathbf{f}_{\mathrm{k}}^{\dagger}\mathbf{g}_{\mathrm{k}}]$ represents the cross products of the channel through the subsurface designed for the UE and the channels through subsurfaces designed for other UEs. As in Remark 2, the main feature of (22) is that the summation of large numbers of weighted phases tends to cause considerable cancellation, reducing the importance of this term.
$\mathbb{E}[\mathbf{f}_{\mathrm{k}}^{\dagger}\mathbf{f}_{\mathrm{k}}]=M\beta_{\mathrm{rb}}^{(k)}\beta_{\mathrm{ur}}^{(k)}\bigg[N_{k}\left(\eta_{\mathrm{rb}}^{(k)2}+\zeta_{\mathrm{rb}}^{(k)2}\right)\\ +\sum_{i=1}^{N_{k}}\sum_{j\neq i}\left(\eta_{\mathrm{rb}}^{(k)2}+\zeta_{\mathrm{rb}}^{(k)2}\mathbf{A}_{i,j}\right)F_{R}\bigg],$ (23)
where
$\mathbf{A}=\mathrm{diag}\left(\mathbf{a}_{\mathrm{r}}^{(k)\dagger}\right)\mathbf{R}_{\mathrm{r},k,k}^{(k)}\mathrm{diag}\left(\mathbf{a}_{\mathrm{r}}^{(k)}\right),$ (24)
$F_{R}=\frac{(1-|\rho^{(k)}_{kk}|^{2})^{2}}{1+\kappa_{\mathrm{ur}}^{(k)}}\!\exp{\!\!\left(\!-\frac{2\kappa_{\mathrm{ru}}^{(k)}(1-\mu_{cf})}{1-|\rho^{(k)}_{kk}|^{2}}\!\right)}\!\sum_{m=0}^{\infty}\sum_{n=0}^{m}$
$\times\!\cos(n\phi_{f})\frac{\epsilon_{n}|\rho^{(k)}_{kk}|^{2m-n}}{m!(m-n)!(n!)^{2}}\!\left(\!\frac{\kappa_{\mathrm{ru}}^{(k)}(1+|\rho^{(k)}_{kk}|^{2}-2\mu_{cf}))}{1-|\rho^{(k)}_{kk}|^{2}}\!\!\right)^{\!\!n}\!\!\$
$\times\!\Gamma^{2}\!\!\left(\!\!m\!+\!\frac{3}{2}\!\right)\!{}_{1}F_{1}^{2}\!\!\left(\!m\!+\!\frac{3}{2},n\!+\!1,\frac{\kappa_{\mathrm{ru}}^{(k)}(1+|\rho^{(k)}_{kk}|^{2}-2\mu_{cf}))}{1-|\rho^{(k)}_{kk}|^{2}}\!\right)\!,$ (25)
$\rho_{kk}^{(k)}=\mathbf{R}^{(k)}_{\mathrm{ur},k,k,i,j}$, $\phi_{f}=\angle((1+|\rho_{kk}^{(k)}|^{2})\mu_{cf}\kappa_{\mathrm{ur}}^{(k)}-2\kappa_{\mathrm{ur}}^{(k)}|\rho_{kk}^{(k)}|^{2}+j(1-|\rho_{kk}^{(k)}|^{2})\mu_{sf}\kappa_{\mathrm{ur}}^{(k)})$, $\mu_{cf}\!=\!\rho^{(k)}_{kk}\!\cos\!\left(\!\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}\!-\!\angle\mathbf{a}_{\mathrm{ur},k,j}^{(k)}\!\right)$ and $\mu_{sf}\!=\!\rho^{(k)}_{kk}\!\sin\!\left(\!\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}\!-\!\angle\mathbf{a}_{\mathrm{ur},k,j}^{(k)}\!\right)$.
Remark 4
$\mathbb{E}[\mathbf{f}_{\mathrm{k}}^{\dagger}\mathbf{f}_{\mathrm{k}}]$ represents the gain of the channel through the subsurface designed for UE $k$. It includes Gaussian hypergeometric functions, which increase monotonically with the final argument. The final argument increases with the spatial correlation between subsurface elements, and this increase occurs when subsurface elements are placed closer together. Hence, placing the subsurface elements close together actually increases the mean SNR.
$\mathbb{E}[\mathbf{g}_{\mathrm{k}}^{\dagger}\mathbf{g}_{\mathrm{k}}]=\frac{M\beta_{\mathrm{rb}}^{(k)}\beta_{\mathrm{ur}}^{(k)}}{\left(\kappa_{\mathrm{rb}}^{(k)}+1\right)\!\!\left(\kappa_{\mathrm{ur}}^{(k)}+1\right)}\Bigg[\sum_{s=t\neq k}\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}\,G_{R}$
$\quad\times\mathrm{e}^{-j\angle\mathbf{a}_{\mathrm{r},s,i}^{(s)}}\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{r},s,j}^{(s)}}\Big(\kappa_{\mathrm{ur}}^{(k)}\mathbf{a}_{\mathrm{ur},s,i}^{(k)*}\mathbf{a}_{\mathrm{ur},s,j}^{(k)}+\rho^{(k)}_{ss}\Big)\Big(\kappa_{\mathrm{rb}}^{(k)}$
$\quad\times\mathbf{a}_{\mathrm{r},s,i}^{(k)}\mathbf{a}_{\mathrm{r},s,j}^{(k)*}+r^{(k)}_{ss}\Big)\!+\!\!\sum_{s\neq t\neq k}\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}(\mathbf{C}^{*}\!(s)\!)_{i,i}(\mathbf{C}(t)\!)_{j,j}$
$\quad\times\!\Big(\!\kappa_{\mathrm{ur}}^{(k)}\mathbf{a}_{\mathrm{ur},s,i}^{(k)*}\mathbf{a}_{\mathrm{ur},t,j}^{(k)}\!+\!\rho^{(k)}_{st}\!\Big)\!\Big(\!\kappa_{\mathrm{rb}}^{(k)}\mathbf{a}_{\mathrm{r},s,i}^{(k)}\mathbf{a}_{\mathrm{r},t,j}^{(k)*}\!+\!r^{(k)}_{st}\!\Big)\!\Bigg$]\!, (26)
where
$G_{R}=\!\frac{2\!\left(\!1\!+\!\kappa_{\mathrm{ur}}^{(s)}\!\right)^{\!\!2}\!\mathrm{e}^{j\!\left(\!\angle(\!1-\mu_{c}+j\mu_{s}\!)+\!\angle\mathbf{a}^{(s)}_{\mathrm{ur},s,j}\!\right)}\!}{\!\left(\!1\!-\!\rho_{ss}^{(s)2}\!\right)\!\mathrm{e}^{j\!\left(\!\angle(\!1-\mu_{c}-j\mu_{s}\!)+\!\angle\mathbf{a}^{(s)}_{\mathrm{ur},s,i}\!\right)}\!}\!\exp\!\!\left(\!\!\frac{-2\kappa_{\mathrm{ur}}^{(s)}\!(\!1\!-\!\mu_{c})}{1\!-\!\rho_{ss}^{(s)2}}\!\!\right)\!$
$\times\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}\frac{\epsilon_{n}\kappa_{0}^{2m+n}}{2^{m+n}m!\Gamma(m\!+\!n\!+\!1)}\left(\mathcal{F}(3)\mathrm{e}^{jn\phi}\!+\!\mathcal{F}(1)\mathrm{e}^{-jn\phi}\right)\!,$ (27)
$\mathcal{F}(x)\!=\!\frac{\Gamma^{2}\!\left(n\!+\!m\!+\!\frac{x}{2}\right)\!\!\left(\!\zeta\sqrt{\kappa_{\mathrm{ur}}^{(s)}(1\!+\!|\rho_{ss}^{(s)}|^{2}\!-\!2\mu_{c}}\right)^{\!\!\!2n-4+2x}}{4^{n-1+x}\,\Gamma^{2}(n-1+x)\left(\frac{1+\kappa^{(s)}_{\mathrm{ur}}}{1-\rho_{ss}^{(s)2}}\right)^{2m+2n+x}}$
$\!\!\!\times\!\!{}_{1}F_{1}^{2}\!\!\left(\!\!n\!+\!m\!+\!\frac{x}{2},\!n\!-\!1\!+\!x,\!\frac{\zeta^{2}\!\kappa_{\mathrm{ur}}^{(s)}\!\!\left(\!1\!+\!|\rho_{ss}^{(s)}|^{2}\!-\!2\mu_{c}\!\right)\!\!\!\left(\!1\!-\!\rho_{ss}^{(s)2\!}\right)}{4\left(1\!+\!\kappa_{\mathrm{ur}}^{(s)}\right)}\!\!\right)\!\!,$ (28)
and $\mathbf{C}(x)$ is given in (21), $\rho^{(k)}_{ss}\!=\!\mathbf{R}^{(k)}_{\mathrm{ur},s,s,i,j}$, $r_{ss}^{(k)}=\mathbf{R}^{(k)}_{\mathrm{r},s,s,i,j}$, $\rho^{(k)}_{st}=\mathbf{R}^{(k)}_{\mathrm{ur},s,t,i,j}$, $r^{(k)}_{st}=\mathbf{R}^{(k)}_{\mathrm{r},s,t,i,j}$, $\rho^{(s)}_{ss}=\mathbf{R}^{(s)}_{\mathrm{ur},s,s,i,j}$, $\mu_{c}\!=\!\rho^{(s)}_{ss}\!\cos\!\left(\!\angle\mathbf{a}_{\mathrm{ur},s,j}^{(s)}\!-\!\angle\mathbf{a}_{\mathrm{ur},s,i}^{(s)}\!\right)$, $\mu_{s}\!=\!\rho^{(s)}_{ss}\!\sin\!\left(\!\angle\mathbf{a}_{\mathrm{ur},s,j}^{(s)}\!-\!\angle\mathbf{a}_{\mathrm{ur},s,i}^{(s)}\!\right)$, $\kappa_{0}\!=\!\frac{2|\rho^{(s)}_{ss}|}{1\!-\!|\rho^{(s)}_{ss}|^{2}}\!\!\left(\!1\!+\!\kappa_{\mathrm{ur}}^{(s)}\!\right)$, $\zeta=\frac{2\sqrt{1+\kappa_{\mathrm{ur}}^{(s)}}}{1-|\rho^{(s)}_{ss}|^{2}}$, $\phi\!=\!\angle(1\!-\!\mu_{c}\!+\!j\mu_{s})\!+\!\!\angle\mathbf{a}_{\mathrm{ur},s,j}^{(s)}\!\!-\!\!\angle(1\!-\!\mu_{c}\!-\!j\mu_{s})\!-\!\!\angle\mathbf{a}_{\mathrm{ur},s,i}^{(s)}\!+\!\angle\rho^{(s)}_{ss}\!$, $\epsilon_{0}=1$, and $\epsilon_{n}=2$ when $n\geq 1$ from [26].
Remark 5
$\mathbb{E}[\mathbf{g}_{\mathrm{k}}^{\dagger}\mathbf{g}_{\mathrm{k}}]$ represents the sum of the gains of the channels through subsurfaces designed for other UEs. Increased correlation between elements of any subsurface is still beneficial to UE $k$, again due to the presence of Gaussian hypergeometric functions that increase monotonically with the spatial correlation of RIS elements. Therefore, it is beneficial to all UEs if all RIS elements are closely collocated.
Remark 6
As the terms in (23) and (26) are quadratic, and thus typically larger than the cross product terms, a change in spatial correlation at the RIS is more influential to the overall SNR than a change in spatial correlation at the BS.
Combining (18), (19), (20), (22), (23) and (26) with (17) gives the complete mean SNR.
IV-B Special Case 1: Mean SNR for Correlated Ricean $\mathbf{H}_{\mathrm{rb}}$ with Correlated Rayleigh UE Channels
Now consider the $\mathbb{E}[\mathrm{SNR}_{k}]$ subcase where $\mathbf{H}_{\mathrm{rb}}$ is correlated Ricean and the UE channels are correlated Rayleigh.
Assuming $\mathbf{h}^{(k)}_{\mathrm{d}}=\sqrt{\beta_{\mathrm{d}}^{(k)}}\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{SC})}$ and $\mathbf{h}^{(k)}_{\mathrm{ur}}=\sqrt{\beta_{\mathrm{ur}}^{(k)}}\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{SC})}$ are correlated Rayleigh channels, then $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)}]=\mathbb{E}[\mathbf{h}_{\mathrm{ur}}^{(k)}]=\mathbf{0}$. As $\mathbf{h}^{(k)}_{\mathrm{d}}$ and $\mathbf{h}^{(k)}_{\mathrm{ur}}$ are independent, and $\mathbf{h}_{\mathrm{ur}}^{(k)}$ and $\mathbf{h}_{\mathrm{ur}}^{(s)}$ are independent for $s\neq k$ , $\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{g}_{k}]$ and $\mathbb{E}[\mathbf{h}^{(k)\dagger}_{\mathrm{d}}\mathbf{g}_{\mathrm{k}}]$ contain zero-mean terms. Therefore, (17) simplifies to
$\mathbb{E}[\mathrm{SNR}_{k}]\!=\!\frac{E_{s}}{\sigma^{2}}\mathbb{E}\Big[\mathbf{h}^{(k)\dagger}_{\mathrm{d}}\mathbf{h}^{(k)}_{\mathrm{d}}\!+2\Re(\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{f}_{k})\!+\!\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}\!+\!\mathbf{g}_{k}^{\dagger}\mathbf{g}_{k}\Big]\!.$ (29)
The result for $\mathbb{E}[\mathbf{h}^{(k)\dagger}_{\mathrm{d}}\mathbf{h}^{(k)}_{\mathrm{d}}$] is the same as (18), while (19), (23) and (26) no longer rely on UE link K-factors and can be simplified to
[…]
$\mathbb{E}[\mathbf{f}_{\mathrm{k}}^{\dagger}\mathbf{f}_{\mathrm{k}}]=M\beta_{\mathrm{rb}}^{(k)}\beta_{\mathrm{ur}}^{(k)}\bigg[N_{k}\left(\eta_{\mathrm{rb}}^{(k)2}+\zeta_{\mathrm{rb}}^{(k)2}\right)+\frac{\pi}{4}\sum_{i=1}^{N_{k}}\sum_{j\neq i}\\ \times\!\left(\!\eta_{\mathrm{rb}}^{(k)2}\!\!+\!\zeta_{\mathrm{rb}}^{(k)2}\!\mathbf{A}_{i,j}\!\right)\!{}_{2}F_{1}\!\left(\frac{-1}{2},\frac{-1}{2},1,|\rho_{kk}^{(k)}|^{2}\!\right)\!\!\bigg],$ (31)
$\mathbb{E}[\mathbf{g}_{\mathrm{k}}^{\dagger}\mathbf{g}_{\mathrm{k}}]=M\beta_{\mathrm{rb}}^{(k)}\beta_{\mathrm{ur}}^{(k)}\frac{\pi}{4}\sum_{s\neq k}\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}\Big(\eta_{\mathrm{rb}}^{(k)2}\mathbf{a}_{\mathrm{r},s,i}^{(k)}\mathbf{a}_{\mathrm{r},s,j}^{(k)*}\\ +\zeta_{\mathrm{rb}}^{(k)2}r_{ss}^{(k)}\!\Big)\rho_{ss}^{(k)}\rho_{ss}^{(s)}\mathbf{a}_{\mathrm{r},s,i}^{(s)*}\mathbf{a}_{\mathrm{r},s,j}^{(s)}{}_{2}F_{1}\left(\frac{1}{2},\frac{1}{2},2,|\rho_{ss}^{(s)}|^{2}\right)\!.$ (32)
Substituting (18), (30), (31) and (32) into (29) gives the mean SNR.
Remark 7
The contribution in (30) is enhanced by alignment between the scattering in the direct link and LoS component of the RIS link. Specifically, alignment between the leading eigenvectors of $\mathbf{R}_{\mathrm{d}}^{(k)}$ and $\mathbf{a}_{\mathrm{b}}^{(k)}$ increases the SNR. In (31), increasing channel power enhances the dominant first term. The second summation term tends to increase with spatial correlation due to the monotonically increasing Gaussian hypergeometric function, while the summation over $\mathbf{A}_{i,j}$ terms is less significant due to the cancellation inherent in summing over phases. The same pattern occurs in (32) with increasing correlation boosting the positive $i=j$ component and $i\neq j$ terms tending to cancel out.
IV-C Special Case 2: Mean SNR for LoS $\mathbf{H}_{\mathrm{rb}}$ with Correlated Rayleigh UE Channels
Now consider the further simplification of $\mathbb{E}[\mathrm{SNR}_{k}]$ for LoS $\mathbf{H}_{\mathrm{rb}}$ and correlated Rayleigh UE links.
Again assuming $\mathbf{h}^{(k)}_{\mathrm{d}}=\sqrt{\beta_{\mathrm{d}}^{(k)}}\tilde{\mathbf{h}}_{\mathrm{d}}^{(k,\mathrm{SC})}$ and $\mathbf{h}^{(k)}_{\mathrm{ur}}=\sqrt{\beta_{\mathrm{ur}}^{(k)}}\tilde{\mathbf{h}}_{\mathrm{ur}}^{(k,\mathrm{SC})}$, let $\mathbf{H}_{\mathrm{rb}}^{(k)}=\sqrt{\beta_{\mathrm{rb}}^{(k)}}\tilde{\mathbf{H}}_{\mathrm{rb}}^{(k,\mathrm{LoS})}$. Due to the independence of $\mathbf{h}^{(k)}_{\mathrm{d}}$ and $\mathbf{h}^{(k)}_{\mathrm{ur}}$, the $\mathbb{E}[\mathrm{SNR}_{k}]$ expression is equivalent to (29). $\mathbb{E}[\mathbf{h}^{(k)\dagger}_{\mathrm{d}}\mathbf{h}^{(k)}_{\mathrm{d}}$] is again equivalent to (18), while the remaining terms are no longer dependent on the RIS-BS link K-factor and can be further simplified to
[…]
$\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}]\!=\!M\!\beta_{\mathrm{rb}}^{(k)}\!\beta_{\mathrm{ur}}^{(k)}\!\bigg(\!\!N_{k}\!+\!\frac{\pi}{4}\!\sum\limits_{i=1}^{N_{k}}\!\sum\limits_{j\neq i}{}_{2}F_{1}\!\!\left(\!\!\frac{-1}{2},\frac{-1}{2};1;|\rho_{kk}^{(k)}|^{2}\!\right)\!\!\!\bigg),$ (34)
$\mathbb{E}[\mathbf{g}_{\mathrm{k}}^{\dagger}\mathbf{g}_{\mathrm{k}}]=M\beta_{\mathrm{rb}}^{(k)}\beta_{\mathrm{ur}}^{(k)}\frac{\pi}{4}\sum_{s\neq k}\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}\rho_{ss}^{(k)}\rho_{ss}^{(s)}\mathbf{a}_{\mathrm{r},s,i}^{(k)}\mathbf{a}_{\mathrm{r},s,j}^{(k)*}\\ \times\mathbf{a}_{\mathrm{r},s,i}^{(s)*}\mathbf{a}_{\mathrm{r},s,j}^{(s)}{}_{2}F_{1}\left(\frac{1}{2},\frac{1}{2},2,|\rho_{ss}^{(s)}|^{2}\right)\!.$ (35)
Substituting (18), (33), (34) and (35) into (29) gives the mean SNR, matching those results we first reported in [16].
Remark 8
As in Remark 7, alignment between $\mathbf{R}_{\mathrm{d}}^{(k)}$ and $\mathbf{a}_{\mathrm{b}}^{(k)}$ and increased spatial correlation tends to increase the SNR components in (33)-(35).
V Numerical Results
Numerical results were generated to verify the analysis above and explore the performance of the subsurface design. The channel gain values are selected based on the distance based path loss model detailed in [19], where
$\beta=C_{0}\left({d}/{D_{0}}\right)^{-\alpha},$ (36)
$D_{0}$ is the reference distance of 1 m, $C_{0}$ is the path loss at $D_{0}$ (-30 dB), $d$ is the link distance in metres and $\alpha$ is the path loss exponent ($\alpha_{\mathrm{d}}=\alpha_{\mathrm{rb,NLoS}}=3.5$, $\alpha_{\mathrm{rb,LoS}}=2$ and $\alpha_{\mathrm{ru}}=2.8$).
It is assumed that the RIS is fixed at 40 m away from the BS, at an angle of $\frac{\pi}{4}$ rad. Users are then dropped in a straight corridor of width $2w$ between BS and RIS, at a maximum distance of $r$ from the RIS with an exclusion zone of 1 m around the RIS. This region ensures that the path from the UE through the RIS significantly contributes to the total channel, and the effects of the RIS can be clearly seen. User clusters are considered to investigate the effect of people naturally congregating. Three types of user grouping were used in simulations. As seen in Fig. 2, layout A places all $K$ UEs in one cluster, where they are all in the same area spaced 1 m apart and layout B creates 2 clusters of $\frac{K}{2}$ UEs, with users spaced 1 m apart. Cluster locations are drawn from a uniform distribution. Layout C places each of the $K$ UEs randomly within the corridor, again using a uniform distribution. To ensure that results are general and not tied to a specific UE location, $10^{4}$ random user drops are generated per layout, each with $10^{6}$ replicates.
Figure 2: User drop configurations for one cluster (a), two clusters (b) and random (c).
For simulation purposes, we use the Rayleigh fading correlation model proposed in [27], where
$\mathbf{R}_{n,m}=\mathrm{sinc}\left(2\,d_{mn}\right),\quad n,m=1,\dots,L,$ (37)
$\mathbf{R}\in\{\mathbf{R}_{\mathrm{d}},\mathbf{R}_{\mathrm{b}},\mathbf{R}_{\mathrm{r}},\mathbf{R}_{\mathrm{ur}}\}$, $d_{mn}$ is the Euclidean distance between BS antennas/ RIS elements $m$ and $n$, measured in wavelength units, and $L$ is the number of BS antennas/ RIS elements.
The steering vectors $\mathbf{a}_{\mathrm{d}}$, $\mathbf{a}_{\mathrm{b}}$, $\mathbf{a}_{\mathrm{r}}$ and $\mathbf{a}_{\mathrm{ur}}$ correspond to the vertical uniform linear array (VURA) model and are given in [28]. Elements are arranged in a VURA at intervals of $d_{b}$ and $d_{r}$ wavelengths at the BS and RIS, respectively. $M_{x}$ and $N_{x}$ are the number of BS and RIS elements per row and $M_{z}$ and $N_{z}$ are the number of BS and RIS elements per column, such that $M=M_{x}M_{z}$ and $N=N_{x}N_{z}$. $\theta_{\mathrm{D}}$ and $\phi_{\mathrm{D}}$ are the elevation and azimuth angles-of-departure (AoDs) at the RIS and $\theta_{\mathrm{A}}$ and $\phi_{\mathrm{A}}$ are the corresponding elevation and azimuth angles-of-arrival (AoAs) at the BS. We assume the RIS is on a $\frac{\pi}{4}$ rad angle with respect to the BS, so $\phi_{\mathrm{D}}=\frac{5\pi}{4}$ rad and $\phi_{\mathrm{A}}=\frac{\pi}{4}$ rad. We also assume both are at the same height, so $\theta_{\mathrm{D}}=\theta_{\mathrm{A}}=\frac{\pi}{2}$ rad.
Without loss of generality we assume $\sigma^{2}=1$ and $E_{s}$ is selected in Sec V-A so that the mean output SNR for one UE is 5 dB in a system using the SD where $N=128$, $M=16$, $d_{r}=0.1$, $d_{b}=0.5$ and $\kappa_{d}=\kappa_{rb}=\kappa_{ur}=1$. In Sec. V-B, we select $E_{s}$ so that the mean output SNR for one UE is 5 dB in a system using the SD where $N=128$, $M=16$, $d_{r}=0.1$, $d_{b}=0.5$, $\kappa_{\mathrm{ur}}=\kappa_{\mathrm{d}}=1$ and the RIS-BS channel is purely LoS, an important special case. These parameter values and definitions do not change throughout the results, unless specified otherwise.
V-A Mean SNR Comparison of Subsurface Methods in Sec. III
Fig. 3 verifies the analytical results in Sec. IV-A by showing the equivalence of $\mathbb{E}[\mathrm{SNR}]$ and the simulated mean SNR for the SD, and then compares this with the simulated mean SNR for the ISD. Two versions of the ISD were implemented. The first iterates through all subsurfaces once, as shown in Sec. III-C. The second iterates through all subsurfaces and calculates the SNR of the system. It then repeats this process multiple times until the SNR of the system has converged, i.e. it increases by less than a specified tolerance from one iteration to the next. For these simulations, the tolerance was specified to be $10^{-4}$. This method is referred to as the CISD. The array parameters for Fig. 3 are $M=16$, $M_{x}=4$, $d_{b}=0.5$ and all K-factors are 1.
Figure 3: Comparison of mean SNR for a randomly selected UE for different subsurface phase selection methods while varying the number of RIS elements.
It can be seen that the simulated mean SNR for the method outlined in Sec. III-A is virtually indistinguishable from the $\mathbb{E}[\mathrm{SNR}]$ derived in Sec. IV-A, verifying that result. As expected, the mean SNR for all 3 methods increases with $N$, however the rate of improvement decreases as $N$ increases. This is expected, as it is well known that the RIS SNR can grow quadratically with $N$ [18] so that the dB relationship is logarithmic. The quadratic behavior can be seen in (23), (31), and (34) where the $\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}]$ term contains the sum of $N_{k}(N_{k}-1)$ positive terms, giving a contribution of $\mathcal{O}(N_{k}^{2})$. The ISD and converged ISD both provide more benefit at higher $N$. At high $N$, more improvement can be made by iterating, as there are more elements to vary and hence more opportunities to improve the performance of the system. This improvement is more pronounced when the correlation between RIS elements is lower. Figure 3 shows that the gap between the mean SNR of the SD and CISD is larger for RIS element spacings of $d_{r}=0.5$ than for $d_{r}=0.1$. This may be due to the fact that there is more variability in the previously set subsurfaces creating the fixed channels for the iterative design. Hence, reduced correlation leads to greater variation in the fixed channels, which the iterative design can exploit. Additionally, when $N$ is kept constant, the average SNR per UE decreases as the number of UEs increases. This reduction arises because the fixed set of RIS elements must be divided between more UEs, resulting in fewer elements being allocated to each UE.
A key design choice for the ISD is the order in which subsurfaces are set. The standard ISD in Sec. III-C sets the subsurface of weaker UEs first. This allows the strongest UEs to benefit from additional channel information, and leads to higher average performance. However, this does not lead to fair performance across all UEs. Table I compares the ISD to two other versions. The first is a reverse version, where strongest UEs are set first, allowing the weakest UEs to benefit from extra information. The second is a random version, where the order in which the UEs are set is random. This version acts as a benchmark and allows insights into the respective benefits provided by the ISD and reverse ISD. The mean SNR of the strongest UE (UE 1) to the weakest UE (UE 4) for all three methods is shown for $N=128$ and $d_{r}=0.5$.
TABLE I: Mean SNR ($\mathrm{dB}$) for each UE using the ISD and the reverse ISD.
UE 1 UE 2 UE 3 UE 4
ISD 15.12 7.87 3.35 0.30
Reverse ISD 13.38 7.25 3.44 0.93
Random ISD 14.42 7.51 3.35 0.63
The reverse ISD offers fairer performance at the expense of a rate drop for the stronger UEs and a lower average rate. The random ISD provides a balance between the two, with less of a rate drop than the reverse ISD, but also lower fairness. Thus, the choice of subsurface ordering should depend on whether fairness or average rate is the priority.
The increase in performance provided by the ISD and CISD also leads to an increase in required CSI. Both methods require the same amount of CSI as each needs to estimate $N$ channels between the RIS and the UE, as opposed to the $\frac{N}{K}$ required by the SD. This is due to the need to measure the impact of each subsurface on all users, rather than on just the user it is designed for. Gathering CSI is expected to be challenging for RIS systems, so the benefit from an iterative method would need to be weighed up against this additional CSI requirement.
Each iteration of the ISD is very simple. Additional complex multiplies are needed to compute channel norms to order subsurfaces by impact and to recalculate fixed channel components. For the CISD, the SNR also needs to be computed each iteration. Table II compares the complexity of the three subsurface methods if $n$ iterations are required.
TABLE II: Complexity comparison of the subsurface methods.
SD ISD CISD
Multiplies KM+NKM\!+\!N $NK\!+\!n\!\Big(\!\frac{(K\!-\!1)!MN}{K}$ $\times\!\!\left(\!\frac{N}{K}\!+\!1\!\right)\!+\!KM\!+\!N\!\Big)$ $NK\!+\!n\!\Big(\!\frac{(K\!-\!1)!MN}{K}$ $\times\!\!\left(\!\frac{N}{K}\!+\!1\!\right)\!+\!KM\!(2\!+\!N(N\!+\!1)\!)\!+\!N\Big)$
Divides $2N+K$ $n(2N+K)$ $n(2N+K)$
Mathematical Functions None 1 sort($\cdot$)
1 index($\cdot$) 1 sort($\cdot$)
1 index($\cdot$)
Array
Functions None 1 sort($\cdot$)
1 index($\cdot$) 1 sort($\cdot$)
1 index($\cdot$)
The average number of iterations required for convergence for each $N$ is shown in Fig. 4. The largest mean number of iterations for an $M=16$, $N=128$ system occurs when $d_{r}=0.5$ and is 6.60. In this scenario, when the maximum number of iterations is fixed at 7, the resulting SNR averages to be 99.8% of the fully converged SNR. Hence, restricting the number of iterations to a fixed amount has little impact on performance.
Figure 4: Mean number of iterations of the ISD required for convergence.
V-B Mean Sum Rate of SD versus Non-Subsurface Methods
Many existing MU RIS phase selection methods, such as those in [4, 8, 7, 10], focus on achieving near optimal performance. Necessarily, they have challenging computational complexity requirements. The lowest complexity MU method we are aware of is that proposed in [9]. While other methods may offer better performance, the design in [9] is more likely to be considered for practical implementation. This method approximately minimizes the total mean-squared error (TMSE) of a minimum mean-squared error (MMSE) receiver. All users are placed in one band, so each user has $K$ times the bandwidth (BW) when compared to users served by the SD. This discrepancy in available BW per UE would exist between the SD and any MU approach where all of the UEs are in a single frequency band. We believe it is important to compare our method to this type of scenario, as this is the most common type of MU RIS design in the literature. Also, considering a key benefit of the SD is a significant complexity reduction, a low-complexity MU approach was considered to be a fair comparison. To show the low complexity nature of the TMSE method and the significant complexity reduction afforded by the SD, the number and complexity of matrix operations required for both methods is compared to the low-complexity near-optimal approach in [10] in Table III. The approach in [10] jointly optimizes the RIS phase shifts, precoder and combiner for a single UE system to minimize the MSE. For a fair comparison with the SD and TMSE approaches, the number of operations for the single UE case are scaled by $K$, representing an OFDM-type system with no IUI.
TABLE III: Complexity comparison of the SD, the TMSE method in [9] and the MSE method in [10].
SD TMSE [9] MSE [10]
Complex Multiplicative Operations 3N+K(M+1)3N\!+\!K(M\!+\!1) $(N\!+\!M)(K\!+\!K^{2})\!+\!K(3\!+\!3K+\!2K^{2})$ nK(2N(M+1)+M+3)nK\!(2N(M\!+\!1)\!+\!M\!+\!3)
Matrix Functions None 2 $K\!\times\!K$ inverses
1 $K\!\times\!K$ eigendecomposition 1 $M\times 1$ eigendecomposition
Assuming that $N$ is the largest variable, the SD offers a multiplicative reduction of a factor of $K+K^{2}$ over the TMSE method, and of $nKM$ over the MSE approach. For a system where $N=128$, $M=32$, $K=4$ and $n=40$ (as suggested in [10]), the SD, TMSE method and MSE method require 516, 3388 and 1356960 multiplicative operations, respectively, highlighting the significant computational savings provided by the SD. Additionally, the SD does not require matrix functions, unlike the others. Results in [16] also showed that the SD provides fairer results according to Jain’s fairness test, uses a much simpler receiver type (MF vs MMSE) and reduces CSI requirements by a factor of $K$ when compared to the TMSE approach. Hence, the SD approach offers substantial complexity savings even compared to the simplest RIS designs.
In this section, the mean sum rates for the SD and TMSE methods are compared with the mean sum rate for randomly selected phases. The inclusion of random phases creates a lower bound to compare both methods against. For the random approach, the phase shift of each element is generated from a uniform distribution between 0 and $2\pi$. It is assumed that all users are in one frequency band like the TMSE example. We investigate the impact of important system features, such as K-factor levels, correlation, system size and UE locations.
V-B1 K-Factors and Clustering
Figs. 5, 6 and 7 compare the mean sum rate for the SD with that of random phases and the TMSE design for varying levels of the UE-BS, RIS-BS and UE-RIS K-factor, respectively. $\eta$, as defined in (LABEL:eq:etad), (LABEL:eq:etarb) and (LABEL:eq:etaur), rather than K-factor itself is plotted so the proportion of the link that is LoS can be easily observed.
The array parameters for Figs. 5, 6 and 7 are M=16M\!=\!16, N=128N\!=\!128, K=4K\!=\!4, $M_{x}\!=\!4$, $N_{x}\!=\!16$, $d_{b}\!=\!0.5$, and $d_{r}\!=\!0.1$. The two fixed K-factors in each plot are set to 1 so that the LoS and NLoS component strengths are equal. Note that RLU represents randomly located UEs and 1 UC and 2 UCs represent one and two clusters of UEs, respectively.
Figure 5: Comparison of mean sum rate for different phase selection methods while varying the UE-BS channel K-factor, and hence $\eta_{\mathrm{d}}$.
Figure 6: Comparison of mean sum rate for different phase selection methods while varying the RIS-BS channel K-factor, and hence $\eta_{\mathrm{rb}}$.
Figure 7: Comparison of mean sum rate for different phase selection methods while varying the UE-RIS channel K-factor, and hence $\eta_{\mathrm{ur}}$.
Remarkably, Figs. 5, 6 and 7 show that despite the significant BW restriction and lower complexity, the SD actually outperforms the TMSE when the K-factor of one of the RIS links is high. A LoS or near-LoS RIS link is a key operating scenario that would be deliberately selected in practice. RIS are often motivated as a method of overcoming propagation challenges in the direct path, and therefore, it is expected that their placement could be selected to ensure a strong LoS path to the BS and/or UEs. In such configurations, a high K-factor is not only realistic but desirable. Thus, the SD excels in a scenario of significant practical relevance and is well positioned for use in deployed RIS systems. While the TMSE method yields higher sum rates at lower K-factors, its complexity scales much more rapidly with system size than that of the SD. Therefore, even for lower K-factor values, the SD has benefits for larger systems.
A higher K-factor is beneficial for the SD but negatively impacts the TMSE design. As the K-factor, and therefore $\eta$ (as defined in (LABEL:eq:etad), (LABEL:eq:etarb) and (LABEL:eq:etaur)) of a channel increases, the rank of that channel decreases until it reaches rank-1 when $\eta=1$. This limits the spatial multiplexing ability of the channel by reducing diversity. Insufficient diversity leads to the inability to separate MU channels. As all users are in separate bands for the SD, a lack of channel diversity is not an issue. The SNR improves as channels can be aligned so that more power is directed towards the user as the rank decreases. However, for the TMSE design where all users are located in one frequency band, the inability to separate channels as the K-factor increases leads to a significant drop in rate. A slight exception to this trend can be seen in Fig. 6 between $\eta_{\mathrm{rb}}=0$ and $\eta_{\mathrm{rb}}=0.3$. Note that the TMSE method assumes a LoS component in $\mathbf{H}_{\mathrm{rb}}$, so increasing the K-factor when $\eta_{\mathrm{rb}}$ is low helps slightly, as the channel becomes more similar to that which the framework was built upon. However, above this threshold, the reduction in diversity dominates and the overall rate reduces. The inability to separate channels also negatively impacts the random design, but this impact is lower as it is already operating at a low rate.
Varying the K-factors for the UE-RIS and RIS-BS channels results in similar mean sum rate behaviour. Data through the RIS must travel through both channels, so changing the diversity of one places limits on the other, and thus on the whole RIS link. This is more prominent when the LoS link is dominant - the mean sum rate for the TMSE method drops dramatically, due to the direct link providing the only diversity. The same can be observed for the random method, but to lesser extent. In comparison, when the UE-BS channel becomes strongly LoS, the mean sum rate for the TMSE design is less impacted. As discussed at the start of Sec. V, UEs are dropped in locations that ensure a strong RIS path, to ensure the effects of the RIS are visible. Therefore, the rate remains high in this case due to the strong RIS link being unaffected.
The impact of placing all four users in a cluster (layout A) is most noticeable in strongly LoS channels, where it improves the rate of the SD and lowers the rate of the TMSE method. Clusters of densely located UEs are expected in realistic scenarios of highly populated areas, such as stadiums or transit hubs. Clustering users reduces channel diversity and negatively impacts the rate of the single frequency TMSE method. A key benefit of the SD is its robustness to clustered users. As it does not rely on spatial multiplexing, clustered users improve its performance due to the similarity of the LoS paths, allowing those subsurfaces designed for other users to provide better signal enhancement. Interestingly, clustering also helps the random method. This is likely due to the fact that when the random phases happen to be beneficial for one user, they benefits all UEs in the cluster.
V-B2 Correlation
Fig. 8 compares the mean sum rate for the SD and TMSE designs for a range of element spacings at the RIS and BS. Elements located closer together result in higher correlation. As these results involve BS element spacings of $d_{b}=0.1\lambda$, the effects of mutual coupling at the BS must be taken into account. Accordingly, for all curves corresponding to $d_{b}=0.1\lambda$, the mutual coupling model in [29] is incorporated. Specifically, the spatial correlation matrices $\mathbf{R}_{\mathrm{b}}$ and $\mathbf{R}_{\mathrm{d}}$ are replaced by their effective counterparts $\mathbf{R}_{\mathrm{b}}^{\mathrm{eff}}=\mathbf{Z}^{\dagger}\mathbf{R}_{\mathrm{b}}\mathbf{Z}$ and $\mathbf{R}_{\mathrm{d}}^{\mathrm{eff}}=\mathbf{Z}^{\dagger}\mathbf{R}_{\mathrm{d}}\mathbf{Z}$, respectively, where $\mathbf{Z}$ is defined in [29, Eq. (9)]. The array parameters are the same as for Figs. 5-7, except for the varied element spacing.
Figure 8: Comparison of mean sum rate for different correlations and phase selection methods while varying the RIS-BS channel K-factor, and hence $\eta_{\mathrm{rb}}$.
For the SD, it is observed that lower element spacing at the RIS increases rate. This verifies the analysis in Sec. IV-A, as (23) monotonically increases with correlation. In contrast, lower element spacing at the BS decreases rate. This also verifies findings in Sec. IV-A, as (19) decreases with correlation. The RIS element spacing is more influential than the BS element spacing to the rate of the SD. As expected, the rate is higher for the scenario where both spacings are $0.1\lambda$ than when both spacings are $0.5\lambda$, even when mutual coupling is considered at the BS. This is due to RIS correlation predominantly affecting the typically larger quadratic $\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}$ term in (23), and BS correlation mostly affecting the typically smaller $\mathbf{h}_{\mathrm{d}}^{\dagger}\mathbf{f}_{k}$ cross product in (19).
The TMSE method is designed for a RIS-BS channel with a strong LoS component, which is assisted by higher correlation and a higher $\eta_{\mathrm{rb}}$. This is evidenced by smaller RIS element spacing increasing the mean sum rate of the TMSE design. However, as $\eta_{\mathrm{rb}}$ also becomes large, there is a substantial performance degradation. While the algorithm is more accurate for this scenario, strongly LoS channels are not ideal to support multiple users, as the MU channels cannot be separated. Also, unlike the SD, the BS element spacing is more influential to the rate of the TMSE design. The TMSE method relies on channel separability to serve multiple users in one band. Decreasing the BS element spacing limits this in both the direct and RIS paths, significantly decreasing the performance of the TMSE method. The SD performs comparatively better when correlation is high at the BS, due to separability not being required for its operation.
V-B3 Array Dimensions
Fig. 9 compares the mean sum rate for 5 combinations of $M$ and $N$ values to investigate the impact of array dimensions. The array parameters are $K=4$, $d_{b}=0.5$, $d_{r}=0.1$, $\kappa_{\mathrm{d}}=1$ and $\kappa_{\mathrm{ur}}=1$.
Figure 9: Comparison of mean sum rate for different array dimensions and phase selection methods while varying the RIS-BS channel K-factor, and hence $\eta_{\mathrm{rb}}$.
In the case of TMSE, it is observed that mean sum rate is influenced significantly more by $M$ than $N$. The largest mean sum rates are for the $M=32$ cases, and then ordered by the size of $N$. Comparatively, the $M=8$ cases give the smallest mean sum rates, with $M=8$, $N=64$ performing the worst. The TMSE method needs sufficient $M$ to multiplex, as $M$ impacts both the direct and RIS links. While a higher $N$ is beneficial, it only affects the RIS link. Therefore, if $N$ is too low or $\eta_{\mathrm{rb}}$ is too high for efficient multiplexing in the RIS path, the direct path could still support it if $M$ is large enough. Table IV shows that the percentage decrease in the mean sum rate from $\eta_{\mathrm{rb}}=0$ to $\eta_{\mathrm{rb}}=1$ is greatest for the TMSE method when $N$ is large and $M$ is small.
TABLE IV: Percentage decrease of TMSE method mean sum rate from $\eta_{\mathrm{rb}}=0$ to $\eta_{\mathrm{rb}}=1$.
$M$ NN Mean sum rate decrease from $\eta_{\mathrm{rb}}=0$ to $1$ (%)
8 256 56.2
32 256 49.6
8 128 45.6
32 128 40.0
8 64 29.4
In contrast, the mean sum rate of the SD increases slightly with $\eta_{\mathrm{rb}}$ and is much more robust in low $M$ scenarios, outperforming the TMSE method for all $M=8$ cases when $\eta_{\mathrm{rb}}\geq 0.9$. $N$ is the more influential parameter for the SD- the highest mean sum rates are from the highest $N$ scenarios, and ordered within these by the highest $M$, which verifies the analysis in Sec. V. As the SD does not rely on multiplexing, $M$ is less important, and higher $N$ leads to more elements designed for each user and more scattering elements. A higher $\eta_{\mathrm{rb}}$ also leads to more power aligned in a specific direction and thus an increased mean sum rate. As percentage increase is a poor metric due to the small magnitude of the change, Table V shows the SD mean sum rate increase from $\eta_{\mathrm{rb}}=0$ to $\eta_{\mathrm{rb}}=1$ for each parameter set.
TABLE V: Mean sum rate increase of SD from $\eta_{\mathrm{rb}}=0$ to $\eta_{\mathrm{rb}}=1$.
$M$ NN Mean sum rate increase from $\eta_{\mathrm{rb}}\!=\!0$ to $1$ (bits/s/Hz)
8 256 2.62
8 128 2.18
8 64 1.62
32 256 1.39
32 128 1.16
The mean sum rate increase for the SD is largest for small $M$ and large $N$. Changing $\eta_{\mathrm{rb}}$ has no effect on the direct link. Therefore, when the direct link is weaker and the RIS link is stronger, a higher percentage of the total channel is being strengthened by the stronger LoS component in the RIS-BS channel. Therefore, the SD method has advantages at high $N$ and $\eta_{\mathrm{rb}}$ and low $M$, where the TMSE method is weakest.
VI Conclusions and Future Work
VI-A Conclusion
In this work, we extended the SD to a higher performing low complexity ISD by setting the subsurfaces sequentially from the weakest user to the strongest so that the users with the largest impact on total SNR benefit from the most prior knowledge. We then proposed the CISD, where the ISD is repeated until the SNR increases by less than a specified threshold. We compared the performance of the SD, ISD and CISD, and found that the iterative methods outperform the SD. The largest benefit from the CISD occurs at high $N$ where there are more elements that can be further improved with increased knowledge. We also derive the exact closed form expression for the mean SNR of the RIS SD where spatially correlated Ricean fading is assumed for all channels, leading to deeper insights into system performance. This is a very general analysis, and thus we show that the expression can be simplified to verify the results in [16] for correlated Rayleigh UE-BS and UE-RIS channels and a LoS RIS-BS channel. We showed that even while having significantly less BW per user, the SD outperformed the much more complex TMSE method for the realistic scenarios when a channel is strongly LoS or users are clustered together.
VI-B Future Work
This work opens up several avenues for future investigation. Firstly, this work focuses on an UL system. The proposed methods are also applicable to DL operation, and under the assumption of perfect CSI, the derived analytical expressions remain structurally unchanged. This assumption is reasonable in the UL, as the required effective channel for RIS phase optimization can be directly inferred from received pilot signals. However, this assumption is less valid in the DL. Relaxing this assumption would require explicitly modeling uncertainty in the channels used for RIS phase design, which would modify the analytical evaluation of the mean SNR. Developing such an analysis constitutes an important direction for future work.
Another important direction for future work is the optimization of the number of RIS elements allocated to each UE. This allocation could be performed either independently of, or jointly with, the RIS phase selection methods proposed in this work. The optimization could be based on a range of criteria, such as each UE’s channel strength, traffic demand, throughput maximization, or fairness objectives. Optimally partitioning the RIS elements among users has the potential to yield significant performance gains and is therefore a promising topic for further research.
Appendix A
Derivation of Mean SNR
We compute the mean SNR by taking the expectation of each term in (17). The first two terms, $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{h}_{\mathrm{d}}^{(k)}]$ and $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)}\mathbf{f}_{k}]$, were derived in [25] with their results listed in (18) and (19). There are four remaining terms to derive.
Term 1: $\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{g}_{k}]$
As the scattered terms are uncorrelated and hence zero mean,
$\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{g}_{k}]$ $=\mathbf{h}_{\mathrm{d}}^{(k,\mathrm{LoS})\dagger}\sum_{s\neq k}\mathbf{H}_{\mathrm{rb},s}^{(k,\mathrm{LoS})}\mathbb{E}[\mathbf{\Phi}_{s}]\mathbf{h}_{\mathrm{ur},s}^{(k,\mathrm{LoS})}.$
Using the definitions for LoS terms in Sec. II-A,
$\mathbb{E}[\mathbf{h}_{\mathrm{d}}^{(k)\dagger}\mathbf{g}_{k}]$ $=\sqrt{\beta^{(k)}_{\mathrm{d}}\beta_{\mathrm{rb}}^{(k)}\beta_{\mathrm{ur}}^{(k)}}\eta^{(k)}_{\mathrm{d}}\eta_{\mathrm{rb}}^{(k)}\eta_{\mathrm{ur}}^{(k)}\mathbf{a}_{\mathrm{d}}^{(k)\dagger}\mathbf{a}_{\mathrm{b}}^{(k)}$
$\qquad\times\sum_{s\neq k}\mathbf{a}_{\mathrm{r},s}^{(k)\dagger}\mathbf{C}(s)\mathbf{a}_{\mathrm{ur},s}^{(k)},$ (38)
where $\mathbf{C}(s)$ is as given in 21 and derived in Appendix B.
Term 2: $\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{g}_{k}]$
Expanding out $\mathbf{f}_{k}$ and $\mathbf{g}_{k}$ gives
$\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{g}_{k}]=\mathbb{E}\bigg[\mathbf{h}_{\mathrm{ur},k}^{(k)\dagger}\mathbf{\Phi}_{k}^{\dagger}\mathbf{H}_{\mathrm{rb},k}^{(k)\dagger}\sum_{s\neq k}\mathbf{H}_{\mathrm{rb},s}^{(k)}\mathbf{\Phi}_{s}\mathbf{h}_{\mathrm{ur},s}^{(k)}\bigg],\\$
$\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{g}_{k}]=\sum_{s\neq k}\mathbb{E}\left[\nu_{k}^{*}\right]\sum_{i=1}^{N_{k}}\sum_{i=1}^{N_{k}}\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|\mathbf{h}_{\mathrm{ur},s,j}^{(k)}\right]\\ \times\mathbb{E}\left[(\mathbf{H}_{\mathrm{rb},k}^{(k)\dagger}\mathbf{H}_{\mathrm{rb},s}^{(k)})_{ij}\right](\mathbf{C}(s))_{j}\mathrm{e}^{-j\angle\mathbf{a}_{\mathrm{r},k,i}^{(k)}}.$ (39)
From (64) and (65) in Appendix B,
$\mathbb{E}[\nu_{k}^{*}]=\frac{\sqrt{\pi\kappa_{\mathrm{d}}^{(k)}}\mathbf{a}_{\mathrm{d}}^{(k)\dagger}\mathbf{a}_{\mathrm{b}}^{(k)}}{2\sqrt{\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{R}_{\mathrm{d}}^{(k)}\mathbf{a}_{\mathrm{b}}^{(k)}}}{}_{1}\!F_{1}\!\left(\frac{1}{2},2,\frac{-|\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{a}_{\mathrm{d}}^{(k)}|^{2}\kappa_{\mathrm{d}}^{(k)}}{\mathbf{a}_{\mathrm{b}}^{(k)\dagger}\mathbf{R}_{\mathrm{d}}^{(k)}\mathbf{a}_{\mathrm{b}}^{(k)}}\!\right)\!\!.$ (40)
As $\mathbf{h}_{\mathrm{ur}}^{(k)}$ is a correlated Ricean vector, we can write $\mathbf{h}_{\mathrm{ur},k,i}^{(k)}=\mathbf{a}_{1}+b\mathbf{u}_{1}$ and $\mathbf{h}_{\mathrm{ur},s,j}^{(k)}=\mathbf{a}_{2}+b_{2}\mathbf{u}_{2}$, where $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$ represent the LoS components, $b_{1}$ and $b_{2}$ represent the scaling of the scattered terms $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$. The Gaussian term of any element can be written as a scaled version of another with noise, so $\mathbf{a}_{2}=\rho\mathbf{u}_{1}+e$ and $\mathbf{h}_{\mathrm{ur},s,j}^{(k)}=\mathbf{a}_{2}+b_{2}(\rho\mathbf{u}_{1}+e)$, where $e$ represents noise and $\rho$ represents the spatial correlation between the elements. Returning to the standard channel definitions in (5) allows us to rewrite $\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|\mathbf{h}_{\mathrm{ur},s,j}^{(k)}\right]$ as
$\mathbb{E}$ $\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|\mathbf{h}_{\mathrm{ur},s,j}^{(k)}\right]\!=\!\sqrt{\frac{\beta_{\mathrm{ur}}^{(k)}\kappa_{\mathrm{ur}}^{(k)}}{(1+\kappa_{\mathrm{ur}}^{(k)})}}\!\!\left(\!\mathbf{a}_{\mathrm{ur},s,j}^{(k)}\!-\!\mathbf{R}_{\mathrm{ur},k,s,i,j}^{(k)}\mathbf{a}_{\mathrm{ur},k,i}^{(k)}\!\right)\!$
$\times\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|\right]+\mathbf{R}_{\mathrm{ur},k,s,i,j}^{(k)}\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|^{2}\mathrm{e}^{j\angle\mathbf{h}_{\mathrm{ur},k,i}^{(k)}}\right]\!.$ (41)
From (4.1) in [30],
$\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|\right]\!\!=\!\!\sqrt{\!\frac{\beta_{\mathrm{ur}}^{(k)}}{(1+\kappa_{\mathrm{ur}}^{(k)})}}\mathrm{e}^{-\kappa_{\mathrm{ur}}^{(k)}}\!\Gamma\!\left(\!\frac{3}{2}\!\right)\!\!{}_{1}F_{1}\!\!\left(\!\frac{3}{2},1,\kappa_{\mathrm{ur}}^{(k)}\!\!\right)\!\!.$ (42)
Using the PDF for a complex random variable (RV) in (3.4) of [30] and letting $r=|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|$, $\theta=\angle\mathbf{h}_{\mathrm{ur},k,i}^{(k)}$ and $a=\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}$,
$\mathbb{E}\left[r^{2}\mathrm{e}^{j\theta}\right]=\int_{0}^{\infty}\!\!\!\!\int_{0}^{2\pi}r^{2}f^{*}(r,\theta)\,d\theta\,dr,$
$=\frac{(1+\kappa_{\mathrm{ur}}^{(k)})\mathrm{e}^{-\kappa_{\mathrm{ur}}^{(k)}}}{\pi\beta_{\mathrm{ur}}^{(k)}}\!\int_{0}^{\infty}\!\!\!r^{3}\mathrm{exp}\left(-\frac{(1+\kappa_{\mathrm{ur}}^{(k)})r^{2}}{\beta_{\mathrm{ur}}^{(k)}}\right)$
$\times\int_{0}^{2\pi}\!\!\!\!\mathrm{e}^{j\theta}\mathrm{exp}\!\!\left(\!\!2\sqrt{\frac{\kappa_{\mathrm{ur}}^{(k)}\!\!\left(1+\kappa_{\mathrm{ur}}^{(k)}\right)}{\beta_{\mathrm{ur}}^{(k)}}}r\cos(\theta\!-\!a)\!\!\right)d\theta dr\!.$ (43)
Using Euler’s formula and the difference of cosines relationship in (1.313.5) and then the integral result in (3.937.2) of [31], the inner integral can be evaluated to give
$\mathbb{E}\left[r^{2}\mathrm{e}^{j\theta}\right]$ $=\frac{2(1+\kappa_{\mathrm{ur}}^{(k)})\mathrm{e}^{-\kappa_{\mathrm{ur}}^{(k)}}}{\beta_{\mathrm{ur}}^{(k)}}\sqrt{\mathrm{exp}(2j\angle a)}\int_{0}^{\infty}\!\!\!r^{3}$
$\times\!\exp\!\left(\!-\frac{(1+\kappa_{\mathrm{ur}}^{(k)})r^{2}}{\beta_{\mathrm{ur}}^{(k)}}\!\right)\!\!I_{1}\!\!\left(\sqrt{\frac{4\kappa_{\mathrm{ur}}^{(k)}(\kappa_{\mathrm{ur}}^{(k)}+1)}{\beta_{\mathrm{ur}}^{(k)}}}\right)\!\!dr\!.$
Rewriting the modified Bessel function as in (8.406.3) and integrating according to (6.631.1) in [31] gives
$\mathbb{E}\left[r^{2}\mathrm{e}^{j\theta}\right]\!=\!\frac{\beta_{\mathrm{ur}}^{(k)}\!\sqrt{\!\kappa_{\mathrm{ur}}^{(k)}}\mathrm{e}^{j\angle a-\kappa_{\mathrm{ur}}^{(k)}}\Gamma\!\left(\frac{5}{2}\right)}{(1+\kappa_{\mathrm{ur}}^{(k)})}{}_{1}F_{1}\!\!\left(\!\frac{5}{2},2,\kappa_{\mathrm{ur}}^{(k)}\!\right)\!.$ (44)
Therefore,
$\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}|\mathbf{h}_{\mathrm{ur},s,j}^{(k)}\right]\!=\!\frac{\beta_{\mathrm{ur}}^{(k)}\!\sqrt{\!\pi\kappa_{\mathrm{ur}}^{(k)}}}{2(1+\kappa_{\mathrm{ur}}^{(k)})}\!\Bigg[\!\Big(\!\mathbf{a}_{\mathrm{ur},s,j}^{(k)}\!-\!\mathbf{R}_{\mathrm{ur},k,s,i,j}^{(k)}\mathbf{a}_{\mathrm{ur},k,i}^{(k)}\!\Big)$
$\!\!\times\!{}_{\!1}{\!F}_{\!1}\!\!\left(\!\frac{3}{2},1,\kappa_{\mathrm{ur}}^{(k)}\!\right)\!\!+\!\frac{3}{2}\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}}\mathbf{R}_{\mathrm{ur},k,s,i,j}^{(k)}\,{}_{\!1}{\!F}_{\!1}\!\!\left(\!\frac{5}{2},\!1,\!\kappa_{\mathrm{ur}}^{(k)}\!\!\right)\!\!\Bigg]\!.\!\!\!$ (45)
Expanding $\mathbf{H}_{\mathrm{rb}}^{(k)}$ into the form seen in (4), taking the expectation of each term and removing zero mean cross products,
$\mathbb{E}\left[(\mathbf{H}_{\mathrm{rb},k}^{(k)\dagger}\mathbf{H}_{\mathrm{rb},s}^{(k)})_{ij}\right]\!$ $=\!\frac{M\beta_{\mathrm{rb}}^{(k)}}{1+\kappa_{\mathrm{rb}}^{(k)}}\Big(\kappa_{\mathrm{rb}}^{(k)}\mathbb{E}\left[\tilde{\mathbf{H}}_{\mathrm{rb},k,i}^{(k,\mathrm{LoS})\dagger}\tilde{\mathbf{H}}_{\mathrm{rb},s,j}^{(k,\mathrm{LoS})}\right]$
$\quad+\mathbb{E}\left[\tilde{\mathbf{H}}_{\mathrm{rb},k,i}^{(k,\mathrm{SC})\dagger}\tilde{\mathbf{H}}_{\mathrm{rb},s,j}^{(k,\mathrm{SC})}\right]\Big),$
$=\!\frac{M\beta_{\mathrm{rb}}^{(k)}}{1+\kappa_{\mathrm{rb}}^{(k)}}\Big(\!\kappa_{\mathrm{rb}}^{(k)}\mathbf{a}_{\mathrm{r},k,i}^{(k)}\mathbf{a}_{\mathrm{r},s,j}^{(k)*}\!+\!\mathbf{R}_{\mathrm{r},k,s,i,j}^{(k)}\!\Big)\!.$ (46)
Finally, $\mathbf{C}(s)$ is defined in (21) and will be derived in Appendix B. Substituting (40), (45), (Appendix A Derivation of Mean SNR) and (21) into (39) gives (22).
Term 3: $\mathbb{E}[\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}]$
Expanding out $\mathbf{f}_{k}$ with $\mathbf{H}_{\mathrm{rb}}^{(k)}$ split into LoS and scattered components (as seen in (4)) and taking the mean value leaves only the squared terms, due to the independence of the cross product terms. Hence,
$\mathbb{E}[$ $\mathbf{f}_{k}^{\dagger}\mathbf{f}_{k}]=\mathbb{E}\left[\mathbf{h}_{\mathrm{ur},k}^{(k)\dagger}\mathbf{\Phi}_{k}^{\dagger}\mathbf{H}_{\mathrm{rb},k}^{(k,\mathrm{LoS})\dagger}\mathbf{H}_{\mathrm{rb},k}^{(k,\mathrm{LoS})}\mathbf{\Phi}_{k}\mathbf{h}_{\mathrm{ur},k}^{(k)}\right]$
$+\mathbb{E}\left[\mathbf{h}_{\mathrm{ur},k}^{(k)\dagger}\mathbf{\Phi}_{k}^{\dagger}\mathbf{H}_{\mathrm{rb},k}^{(k,\mathrm{SC})\dagger}\mathbf{H}_{\mathrm{rb},k}^{(k,\mathrm{SC})}\mathbf{\Phi}_{k}\mathbf{h}_{\mathrm{ur},k}^{(k)}\right]\!=\!T_{1}\!+\!T_{2}.$ (47)
Splitting $T_{1}$ into a sum of its terms, applying the same process as in (Appendix A Derivation of Mean SNR) for $\mathbb{E}[(\mathbf{H}_{\mathrm{rb},k}^{(k,\mathrm{LoS})\dagger}\mathbf{H}_{\mathrm{rb},k}^{(k,\mathrm{LoS}})_{ij}]$, replacing $\mathbf{\Phi}_{k}$ with its expanded form in (12) and simplifying gives
$T_{1}\!=\!M\beta_{\mathrm{rb}}^{(k)}\eta_{\mathrm{rb}}^{(k)2}\!\!\left(\!N_{k}\!+\!\sum_{i=1}^{N_{k}}\sum_{j\neq i}\mathbb{E}\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}||\mathbf{h}_{\mathrm{ur},k,j}^{(k)}|\right]\!\right)\!.$ (48)
From [25],
$\mathbb{E}\!\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}||\mathbf{h}_{\mathrm{ur},k,j}^{(k)}|\right]=\beta_{\mathrm{ur}}^{(k)}F_{R}=\frac{\beta_{\mathrm{ur}}^{(k)}(1-|\rho^{(k)}_{kk}|^{2})^{2}}{(1+\kappa_{\mathrm{ur}}^{(k)})}$
$\!\times\!\exp{\!\left(\!\frac{-2\kappa_{\mathrm{ru}}^{(k)}(1\!-\!\mu_{cf})}{1\!-\!|\rho^{(k)}_{kk}|^{2}}\!\right)}\!\!\sum_{m=0}^{\infty}\sum_{n=0}^{m}\cos(n\phi_{f})\frac{\epsilon_{n}|\rho^{(k)}_{kk}|^{2m-n}}{m!(m\!-\!n)!(n!)^{2}}$
$\!\times\!\left(\!\frac{\kappa_{\mathrm{ru}}^{(k)}(1+|\rho^{(k)}_{kk}|^{2}\!-\!2\mu_{cf}))}{1-|\rho^{(k)}_{kk}|^{2}}\!\!\right)^{\!\!n}\!\!\Gamma^{2}\!\left(\!m\!+\!\frac{3}{2}\!\right)\!{}_{1}F_{1}^{2}\Bigg(\!m\!+\!\frac{3}{2},n\!+\!1,$
$\quad\frac{\kappa_{\mathrm{ru}}^{(k)}(1+|\rho^{(k)}_{kk}|^{2}-2\mu_{cf}))}{1-|\rho^{(k)}_{kk}|^{2}}\!\Bigg)\!,$ (49)
where $\rho_{kk}^{(k)}=\mathbf{R}^{(k)}_{\mathrm{ur},k,k,i,j}$, $\mu_{cf}\!=\!\rho^{(k)}_{kk}\!\cos\!\left(\!\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}\!-\!\angle\mathbf{a}_{\mathrm{ur},k,j}^{(k)}\!\right)$, $\mu_{sf}\!=\!\rho^{(k)}_{kk}\!\sin\!\left(\angle\mathbf{a}_{\mathrm{ur},k,i}^{(k)}\!-\!\angle\mathbf{a}_{\mathrm{ur},k,j}^{(k)}\!\right)$, $\phi_{f}\!=\!\angle(\!(1\!+\!|\rho_{kk}^{(k)}\!|^{2})\mu_{c}\kappa_{\mathrm{ur}}^{(k)}\!\!-\!\!2\kappa_{\mathrm{ur}}^{(k)}\!|\rho_{kk}^{(k)}\!|^{2}\!\!+\!j(1\!-\!|\rho_{kk}^{(k)}\!|^{2})\mu_{s}\kappa_{\mathrm{ur}}^{(k)})$, $\epsilon_{0}=1$, and $\epsilon_{n}=2$ when $n\geq 1$ as defined in [26].
$T_{2}$ can be found using the same process as for $T_{1}$, giving
$T_{2}\!$ $=\!M\beta_{\mathrm{rb}}^{(k)}\zeta_{\mathrm{rb}}^{(k)2}\!\left(\!N_{k}\!+\!\sum_{i=1}^{N_{k}}\sum_{j\neq i}\mathbb{E}\!\left[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}||\mathbf{h}_{\mathrm{ur},k,j}^{(k)}|\right]\!\mathbf{A}_{i,j}\!\right)\!,$ (50)
where
$\mathbf{A}_{i,j}$ $=\left(\mathrm{diag}\big(\mathbf{a}_{\mathrm{r}}^{(k)}\big)^{\dagger}\mathbf{R}_{\mathrm{r},k,k}^{(k)}\mathrm{diag}\big(\mathbf{a}_{\mathrm{r}}^{(k)}\big)\right)_{i,j},$ (51)
and the result from (49) is used for $\mathbb{E}[|\mathbf{h}_{\mathrm{ur},k,i}^{(k)}||\mathbf{h}_{\mathrm{ur},k,j}^{(k)}|]$. Combining (48) and (50) gives (23).
Term 4: $\mathbb{E}[\mathbf{g}_{k}^{\dagger}\mathbf{g}_{k}]$
Splitting $\mathbb{E}[\mathbf{g}_{k}^{\dagger}\mathbf{g}_{k}]$ into two sums,
$\mathbb{E}[\mathbf{g}_{k}^{\dagger}\mathbf{g}_{k}]=\mathbb{E}\Big[\!\!\sum_{s=t\neq k}\!\!\mathbf{h}_{\mathrm{ur},s}^{(k)\dagger}\mathbf{\Phi}_{s}^{\dagger}\mathbf{H}_{\mathrm{rb},s}^{(k)\dagger}\mathbf{H}_{\mathrm{rb},s}^{(k)}\mathbf{\Phi}_{s}\mathbf{h}_{\mathrm{ur},s}^{(k)}\Big]\\ +\mathbb{E}\Big[\!\!\sum_{s\neq t\neq k}\!\!\mathbf{h}_{\mathrm{ur},s}^{(k)\dagger}\mathbf{\Phi}_{s}^{\dagger}\mathbf{H}_{\mathrm{rb},s}^{(k)\dagger}\mathbf{H}_{\mathrm{rb},t}^{(k)}\mathbf{\Phi}_{t}\mathbf{h}_{\mathrm{ur},t}^{(k)}\Big]\!=\!G_{1}\!+\!G_{2}.$ (52)
Rewriting $G_{1}$ as a sum of its terms and taking the expectation of each independent pair of terms gives
$G_{1}=\!\!\!\sum_{s=t\neq k}\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}\mathbb{E}\Big[\big(\mathbf{H}_{\mathrm{rb},s}^{(k)\dagger}\mathbf{H}_{\mathrm{rb},s}^{(k)}\big)_{i,j}\Big]\\ \times\mathbb{E}\big[\mathbf{\Phi}_{\!s,i}^{*}\mathbf{\Phi}_{\!s,j}\big]\mathbb{E}\big[\mathbf{h}_{\mathrm{ur},s,i}^{(k)*}\mathbf{h}_{\mathrm{ur},s,j}^{(k)}\big].$ (53)
Following the procedure used for (Appendix A Derivation of Mean SNR),
$\mathbb{E}\left[\big(\mathbf{H}_{\mathrm{rb},s}^{(k)\dagger}\mathbf{H}_{\mathrm{rb},s}^{(k)}\big)_{i,j}\right]\!=\!\frac{M\!\sqrt{\!\beta_{\mathrm{rb}}^{(k)}}}{1+\kappa_{\mathrm{rb}}^{(k)}}\!\Big(\!\kappa_{\mathrm{rb}}^{(k)}\mathbf{a}_{\mathrm{r},s,i}^{(k)}\mathbf{a}_{\mathrm{r},s,j}^{(k)*}\!+\!\mathbf{R}_{\mathrm{r},s,s,i,j}^{(k)}\!\Big)\!.$ (54)
Expanding out $\mathbf{\Phi}_{s}$ and simplifying the expression gives
$\mathbb{E}[\mathbf{\Phi}_{\!s,i}^{*}\mathbf{\Phi}_{\!s,j}]=\mathrm{e}^{-j\angle\mathbf{a}_{\mathrm{r},s,i}^{(s)}}\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{r},s,j}^{(s)}}\mathbb{E}\!\left[\mathrm{e}^{j\angle\mathbf{h}_{\mathrm{ur},s,i}^{(s)}}\mathrm{e}^{-j\angle\mathbf{h}_{\mathrm{ur},s,j}^{(s)}}\right]\!,$ (55)
where $G_{R}\!=\!\mathbb{E}[\mathrm{e}^{j\angle\mathbf{h}_{\mathrm{ur},s,i}^{(s)}}\mathrm{e}^{-j\angle\mathbf{h}_{\mathrm{ur},s,j}^{(s)}}]$. Appendix C will derive $G_{R}$, with the final result stated in (27). Expanding $\mathbf{h}_{\mathrm{ur}}^{(k)}$ into the form seen in (5), taking the expectation of each term and removing the zero mean cross product terms gives
$\mathbb{E}\big[\mathbf{h}_{\mathrm{ur},s,i}^{(k)*}\mathbf{h}_{\mathrm{ur},s,j}^{(k)}\big]\!=\!\frac{\!\sqrt{\!\beta_{\mathrm{ur}}^{(k)}}}{1+\kappa_{\mathrm{ur}}^{(k)}}\!\Big(\!\kappa_{\mathrm{ur}}^{(k)}\mathbf{a}_{\mathrm{ur},s,i}^{(k)*}\mathbf{a}_{\mathrm{ur},s,j}^{(k)}\!+\!\mathbf{R}_{\mathrm{ur},s,s,i,j}^{(k)}\!\Big)\!.$ (56)
Combining (54), (55) and (56) gives (53). As for $G_{1}$, rewriting $G_{2}$ as a sum of its terms and taking the expectation of each independent group gives
$G_{2}=\sum_{s\neq t\neq k}\!\sum_{i=1}^{N_{k}}\sum_{j=1}^{N_{k}}\mathbb{E}\left[\big(\mathbf{H}_{\mathrm{rb},s}^{(k)\dagger}\mathbf{H}_{\mathrm{rb},t}^{(k)}\big)_{i,j}\right]\\ \times\mathbb{E}\big[\mathbf{\Phi}_{s,i}^{*}\big]\mathbb{E}\big[\mathbf{\Phi}_{t,j}\big]\mathbb{E}\big[\mathbf{h}_{\mathrm{ur},s,i}^{(k)*}\mathbf{h}_{\mathrm{ur},t,j}^{(k)}\big].$ (57)
Again, from (Appendix A Derivation of Mean SNR),
$\left[\big(\mathbf{H}_{\mathrm{rb},s}^{(k)\dagger}\mathbf{H}_{\mathrm{rb},t}^{(k)}\big)_{i,j}\right]\!=\!\frac{M\!\sqrt{\!\beta_{\mathrm{rb}}^{(k)}}}{1+\kappa_{\mathrm{rb}}^{(k)}}\!\Big(\!\kappa_{\mathrm{rb}}^{(k)}\mathbf{a}_{\mathrm{r},s,i}^{(k)}\mathbf{a}_{\mathrm{r},t,j}^{(k)*}\!+\!\mathbf{R}_{\mathrm{r},s,t,i,j}^{(k)}\!\Big)\!.$ (58)
As will be shown in Appendix B,
$(\mathbf{C}^{*}\!(s)\!)_{i,i}\!=\!\mathbb{E}\big[\mathbf{\Phi}^{*}_{s,i}\big]\!=\!\frac{\pi\sqrt{\!\kappa_{\mathrm{d}}^{(s)}\!\kappa_{\mathrm{ur}}^{(s)}}\mathbf{a}_{\mathrm{d}}^{(s)\dagger}\mathbf{a}_{\mathrm{b}}^{(s)}\!}{4\sqrt{\mathbf{a}_{\mathrm{b}}^{(s)\dagger}\mathbf{R}_{\mathrm{d}}^{(s)}\mathbf{a}_{\mathrm{b}}^{(s)}}}\mathrm{e}^{j\left(\!\angle\mathbf{a}_{\mathrm{ur},s,i}^{(s)}\!-\!\angle\mathbf{a}_{\mathrm{r},s,i}^{(s)}\right)}\\ \times{}_{1}F_{1}\!\!\left(\!\frac{1}{2},\!2,\!-\kappa_{\mathrm{ur}}^{(s)}\!\!\right)\!\!{}_{1}F_{1}\!\!\left(\!\!\frac{1}{2},\!2,\!\frac{-|\mathbf{a}_{\mathrm{b}}^{(s)\dagger}\mathbf{a}_{\mathrm{d}}^{(s)}\!|^{2}\kappa_{\mathrm{d}}^{(s)}}{\mathbf{a}_{\mathrm{b}}^{(s)\dagger}\mathbf{R}_{\mathrm{d}}^{(s)}\mathbf{a}_{\mathrm{b}}^{(s)}}\!\right)\!\!,\!$ (59)
$(\mathbf{C}(t)\!)_{j,j}\!=\!\mathbb{E}\big[\mathbf{\Phi}_{t,j}\big]\!=\!\frac{\pi\sqrt{\kappa_{\mathrm{d}}^{(t)}\kappa_{\mathrm{ur}}^{(t)}}\mathbf{a}_{\mathrm{b}}^{(t)\dagger}\mathbf{a}_{\mathrm{d}}^{(t)}}{4\sqrt{\mathbf{a}_{\mathrm{b}}^{(t)\dagger}\mathbf{R}_{\mathrm{d}}^{(t)}\mathbf{a}_{\mathrm{b}}^{(t)}}}\mathrm{e}^{j\left(\angle\mathbf{a}_{\mathrm{r},t,j}^{(t)}\!-\!\angle\mathbf{a}_{\mathrm{ur},t,j}^{(t)}\right)}\!\\ \times{}_{1}\!F_{1}\!\!\left(\!\frac{1}{2},\!2,\!-\kappa_{\mathrm{ur}}^{(t)}\!\!\right)\!\!{}_{1}F_{1}\!\!\left(\!\!\frac{1}{2},\!2,\!\frac{-|\mathbf{a}_{\mathrm{b}}^{(t)\dagger}\mathbf{a}_{\mathrm{d}}^{(t)}|^{2}\kappa_{\mathrm{d}}^{(t)}}{\mathbf{a}_{\mathrm{b}}^{(t)\dagger}\mathbf{R}_{\mathrm{d}}^{(t)}\mathbf{a}_{\mathrm{b}}^{(t)}}\!\right)\!\!.$ (60)
Finally, using the same process as for (56),
$\mathbb{E}\big[\mathbf{h}_{\mathrm{ur},s,i}^{(k)*}\mathbf{h}_{\mathrm{ur},t,j}^{(k)}\big]\!=\!\frac{\!\sqrt{\!\beta_{\mathrm{ur}}^{(k)}}}{1+\kappa_{\mathrm{ur}}^{(k)}}\!\Big(\!\kappa_{\mathrm{ur}}^{(k)}\mathbf{a}_{\mathrm{ur},s,i}^{(k)*}\mathbf{a}_{\mathrm{ur},t,j}^{(k)}\!+\!\mathbf{R}_{\mathrm{ur},s,t,i,j}^{(k)}\!\Big)\!.$ (61)
Therefore combining (58), (59), (60) and (61) gives (57), and combining (53) and (57) gives (52).
Appendix B
Derivation of $\mathbf{C}(x)=\mathbb{E}[\mathbf{\Phi}_{x}]$
Let $\mathbf{C}(x)=\mathbb{E}[\mathbf{\Phi}_{x}]$. From the definition of $\mathbf{\Phi}_{x}$ in (12),
$\mathbb{E}[\mathbf{\Phi}_{x}]=\mathbb{E}[\nu_{x}]\mathrm{diag}\big(\mathrm{e}^{j\angle\mathbf{a}^{(x)}_{\mathrm{r},x}}\big)\mathbb{E}\Big[\mathrm{diag}\big(\mathrm{e}^{-j\angle\mathbf{h}_{\mathrm{ur},x}^{(x)}}\big)\Big].$ (62)
From (4.12) in [30],
$\mathbb{E}[\mathrm{e}^{-j\angle\mathbf{h}_{\mathrm{ur},x}^{(x)}}]\!\!=\!\!\frac{\sqrt{\pi\kappa_{\mathrm{ur}}^{(x)}}\!}{2}\mathrm{diag}\!\left(\!\mathrm{e}^{-j\angle\mathbf{a}_{\mathrm{ur},x}^{(x)}}\!\right)\!\!{}_{1}{\!F}_{\!1}\!\!\left(\!\frac{3}{2},\!2,\!-\kappa_{\mathrm{ur}}^{(x)}\!\!\right)\!\!,\!$ (63)
$\mathbb{E}[\nu_{x}]\!\!=\!\!\frac{\sqrt{\!\pi\kappa_{\mathrm{abhd}}^{(x)}}}{2}\mathrm{diag}\!\left(\!\mathrm{e}^{j\angle\mathbf{a}_{\mathrm{b}}^{(x)\dagger}\mathbf{a}_{\mathrm{d}}^{(x)}}\!\right)\!{}_{1}\!F_{1}\!\!\left(\!\frac{3}{2},2,-\kappa_{\mathrm{abhd}}^{(x)}\!\!\right)\!,$ (64)
where $\kappa_{\mathrm{abhd}}^{(x)}$ is the K-factor of $\mathbf{a}_{\mathrm{b}}^{(x)\dagger}\mathbf{h}_{\mathrm{d}}^{(x)}$ and
$\kappa_{\mathrm{abhd}}^{(x)}=\kappa_{\mathrm{d}}^{(x)}\frac{|\mathbf{a}_{\mathrm{b}}^{(x)\dagger}\mathbf{a}_{\mathrm{d}}^{(x)}|^{2}}{\mathbf{a}_{\mathrm{b}}^{(x)\dagger}\mathbf{R}_{\mathrm{d}}^{(x)}\mathbf{a}_{\mathrm{b}}^{(x)}}.$ (65)
Therefore, combining (62), (63) and (64) gives (21).
Appendix C
Derivation of $G_{R}=\mathbb{E}[\mathrm{e}^{-j\phi_{i}}\mathrm{e}^{j\phi_{j}}]$
Let $G_{R}=\mathbb{E}\left[\mathrm{e}^{j\angle\mathbf{h}_{\mathrm{ur},s,i}^{(s)}}\mathrm{e}^{-j\angle\mathbf{h}_{\mathrm{ur},s,j}^{(s)}}\right]=\mathbb{E}[\mathrm{e}^{-j\phi_{i}}\mathrm{e}^{j\phi_{j}}]$. Using the joint PDF for two complex random variables (RVs) given in (24a) of [26],
$G_{R}\!=\!\!\!\int_{0}^{\infty}\!\!\!\!\int_{0}^{\infty}\!\!\!\!\int_{-\pi}^{\pi}\!\int_{-\pi}^{\pi}\!\!\!\!\mathrm{e}^{\!-j\phi_{i}}\mathrm{e}^{j\phi_{j}}\!f^{*}\!\!\left(r_{\!i},r_{\!j},\mathrm{e}^{j\phi_{i}}\!,\mathrm{e}^{j\phi_{j}}\!\right)\!d{\phi_{i}}d{\phi_{j}}d{r_{\!i}}d{r_{\!j}}.$
Applying the variable transformations $\phi_{i}=\phi_{i}-\angle(1-\mu_{c}-j\mu_{s})-\angle{\mathbf{a}^{(s)}_{\mathrm{ur},s,i}}$ and $\phi_{j}=\phi_{j}-\angle(1-\mu_{c}+j\mu_{s})-\angle{\mathbf{a}^{(s)}_{\mathrm{ur},s,j}}$,
$G_{R}=\!\frac{\left(\!1\!+\!\kappa_{\mathrm{ur}}^{(s)}\!\right)^{\!\!2}\!\mathrm{e}^{j\!\left(\!\angle(\!1-\mu_{c}+j\mu_{s}\!)+\!\angle\mathbf{a}^{(s)}_{\mathrm{ur},s,j}\!\right)}}{\!\pi^{2}\!\!\left(\!1\!-\!\rho_{ss}^{(s)2}\!\right)\!\mathrm{e}^{j\!\left(\!\angle(\!1-\mu_{c}-j\mu_{s}\!)+\!\angle\mathbf{a}^{(s)}_{\mathrm{ur},s,i}\!\right)}\!\!}\!\exp\!\!\left(\!\!\frac{-2\kappa_{\mathrm{ur}}^{(s)}\!(\!1\!-\!\mu_{c})}{1\!-\!\rho_{ss}^{(s)2}}\!\!\right)\!$
$\times\!\!\!\int_{0}^{\infty}\!\!\!\!\int_{0}^{\infty}\!\!\!r_{\!i}r_{\!j}\!\exp\!\!\left(\!\!\frac{-\big(1\!+\!\kappa_{\mathrm{ur}}^{(s)2}\big)}{1-\rho_{ss}^{(s)}}\!\!\left(r_{i}^{2}\!+\!r_{j}^{2}\right)\!\!\right)\!\!\int_{-\pi}^{\pi}\!\int_{-\pi}^{\pi}\!\!\mathrm{e}^{-j\phi_{i}}\mathrm{e}^{j\phi_{\!j}}$
$\times\!\mathrm{e}^{\kappa_{0}r_{i}r_{\!j}\!\cos(\!\phi_{i}\!-\phi_{j}\!-\phi\!)}\mathrm{e}^{\zeta\!\sqrt{\kappa}(r_{i}\cos(\!\phi_{i}\!)+r_{j}\cos(\!\phi_{j}\!)\!)}d\phi_{i}d\phi_{j}dr_{i}dr_{j},$ (66)
where $\rho_{ss}^{(s)}=\mathbf{R}^{(s)}_{\mathrm{ur},s,s,i,j}$, $\mu_{c}=\rho^{(s)}_{ss}\!\cos\!\left(\!\angle\mathbf{a}_{\mathrm{ur},s,j}^{(s)}-\angle\mathbf{a}_{\mathrm{ur},s,i}^{(s)}\!\right)$, $\mu_{s}=\rho^{(s)}_{ss}\!\sin\!\left(\!\angle\mathbf{a}_{\mathrm{ur},s,j}^{(s)}-\angle\mathbf{a}_{\mathrm{ur},s,i}^{(s)}\!\right)$, $\zeta=\frac{2\sqrt{1+\kappa_{\mathrm{ur}}^{(s)}}}{1-|\rho^{(s)}_{ss}|^{2}}$, $\kappa=\kappa_{\mathrm{ur}}^{(s)}(1+|\rho_{ss}^{(s)}|^{2}-2\mu_{c})$, $\kappa_{0}=\frac{2|\rho^{(s)}_{ss}|}{1\!-\!|\rho^{(s)}_{ss}|^{2}}\!\!\left(\!1\!+\!\kappa_{\mathrm{ur}}^{(s)}\!\right)$ and $\phi=\angle(1\!-\!\mu_{c}\!+\!j\mu_{s})+\angle\mathbf{a}_{\mathrm{ur},s,j}^{(s)}-\angle(1\!-\!\mu_{c}\!-\!j\mu_{s})-\angle\mathbf{a}_{\mathrm{ur},s,i}^{(s)}+\angle\rho^{(s)}_{ss}$.
We can now integrate with respect to both $\phi_{i}$ and $\phi_{j}$. Rewriting $\mathrm{e}^{-j\phi_{i}}$ and $\mathrm{e}^{j\phi_{j}}$ using Euler’s formula allows the double integral over the phases to be split into four smaller double integrals. Rewriting $\mathrm{e}^{\kappa_{0}r_{i}r_{j}\cos{(\phi_{i}-\phi_{j}-\phi)}}$ in the form seen in [31, 8.511.4] and the resulting $\cos(n\phi_{j}-n\phi_{i}-n\phi)$ in the form seen in [31, 1.313.5], removing periodic terms with zero-valued integrals, applying the product to sum formula found from [31, 1.313.5] and finally integrating each of the four double integrals using [31, 3.915.2], we get
$G_{R}=\!\frac{2\!\left(\!1\!+\!\kappa_{\mathrm{ur}}^{(s)}\!\right)^{\!\!2}\!\mathrm{e}^{j\!\left(\!\angle(\!1-\mu_{c}+j\mu_{s}\!)+\!\angle\mathbf{a}^{(s)}_{\mathrm{ur},s,j}\!\right)}\!}{\!\left(\!1\!-\!\rho_{ss}^{(s)2}\!\right)\!\mathrm{e}^{j\!\left(\!\angle(\!1-\mu_{c}-j\mu_{s}\!)+\!\angle\mathbf{a}^{(s)}_{\mathrm{ur},s,i}\!\right)}}\!\exp\!\!\left(\!\!\frac{-2\kappa_{\mathrm{ur}}^{(s)}\!(\!1\!-\!\mu_{c})}{1\!-\!\rho_{ss}^{(s)2}}\!\!\right)\!$
$\times\!\!\!\int_{0}^{\infty}\!\!\!\!\int_{0}^{\infty}\!\!\!r_{\!i}r_{\!j}\!\exp\!\!\left(\!\!\frac{-\big(1\!+\!\kappa_{\mathrm{ur}}^{(s)2}\big)}{1-\rho_{ss}^{(s)2}}\!\!\left(r_{i}^{2}\!+\!r_{j}^{2}\right)\!\!\right)\!\!\sum_{n=0}^{\infty}\epsilon_{n}I_{n}(\kappa_{0}r_{i}r_{j})$
$\times\big(J_{n+1}(-j\zeta\sqrt{\kappa}r_{i})J_{n+1}(-j\zeta\sqrt{\kappa}r_{j})\mathrm{e}^{jn\phi}$
$+J_{n-1}(-j\zeta\sqrt{\kappa}r_{i})J_{n-1}(-j\zeta\sqrt{\kappa}r_{j})\mathrm{e}^{-jn\phi}\big)dr_{i}dr_{j},\!\!$ (67)
where $\epsilon_{0}=1$ and $\epsilon_{n}=2$ when $n\geq 1$. To integrate with respect to $r_{i}$ and $r_{j}$, we can rewrite the modified Bessel function $I_{n}(\kappa_{0}r_{i}r_{j})$ as a Bessel function using [31, 8.406.3], expand that function to its series form using [31, 8.402] and integrate each integral using [31, 6.631.1], resulting in (27).